### v1

# Normalized Remainders: Definition, Origins, Backgrounds, Recent Results, and Basic Properties

Abstract.
In this chapter, the author retrospects the ins and outs of the concept of the normalized remainders of the Maclaurin power series expansions of functions, reviews main results on normalized remainders of some elementary functions since 2023, explores the historic backgrounds of the normalized remainders in combinatorics and number theory on the Bernoulli numbers and polynomials, the Stirling numbers and polynomials, and their generalizations by Broder, Carlitz, and Howard, presents several basic properties of normalized remainders, and collects several guesses and problems posed while investigating normalized remainders of several elementary functions.
Key words and phrases:
Maclaurin power series expansion, normalized remainder, definition, background, origin, basic property, guess, problem
1991 Mathematics Subject Classification:
Primary 41A80; Secondary 41A58
1. Definition of the normalized remainder
In this chapter, we use $\mathbb{N}$ and $\mathbb{N}_{0}$ to denote the sets of all positive integers and all nonnegative integers, respectively.
What is the normalized remainders? The concept of normalized remainders in mathematics—particularly in the context of power series expansions—is indeed quite specialized and not widely familiar, even among mathematicians, because it is an innovation that emerged in 2023. In the context of the Maclaurin or Taylor series expansions, a remainder refers to the error term left after truncating the series at a certain degree. A normalized remainder is an essentially modified version of this error term.
Definition 1.1.
Let $f$ be an infinitely differentiable real-valued function defined on an interval $I\subseteq\mathbb{R}$ with $x_{0}\in I$ being an inner point. The informal series expansion of $f$ around the point $x_{0}$ is
$\sum_{n=0}^{\infty}f^{(n)}(x_{0})\frac{(x-x_{0})^{n}}{n!}.$ (1.1)
Its partial sum of the first $n+1$ terms is
$S_{n}(x,x_{0})=\sum_{k=0}^{n}f^{(k)}(x_{0})\frac{(x-x_{0})^{k}}{k!},\quad n\in\mathbb{N}_{0}.$
  1. (1)
    If $f(x_{0})\neq 0$, then we define
    $T_{-1}[f(x),x_{0}]=\frac{f(x)}{f(x_{0})},\quad x\in I.$
  2. (2)
    If $f^{(n+1)}(x_{0})\neq 0$ for $n\in\mathbb{N}_{0}$, then we define
    $T_{n}[f(x),x_{0}]=\begin{dcases}\frac{1}{f^{(n+1)}(x_{0})}\frac{(n+1)!}{(x-x_{0})^{n+1}}[f(x)-S_{n}(x,x_{0})],&x\in I\setminus\{x_{0}\};\\ 1,&x=x_{0}.\end{dcases}$
We call the quantity $T_{n}[f(x),x_{0}]$ the $n$th normalized remainder of the informal series expansion (1.1) of $f$ around $x_{0}$.
When $x_{0}=0$, we simply write $T_{n}[f(x),0]$ as $T_{n}[f(x)]$.
The terminology of the normalized remainder $T_{n}[f(x)]$ was first explicitly appeared in the papers [5, Section 5], [33, Section 5], [30, Remark 7], and [41, 43, 47, 70, 72, 75, 76, 81]. It is also known as normalized tail in [30, Remark 7] and [5, 41, 47, 70, 76, 81] and as Qi’s normalized remainder in the papers [43, 75].
In the work, we will retrospect the origins of the concept of the normalized remainders of the Maclaurin power series expansions of functions, review main results on normalized remainders of some elementary functions since 2023, explore the historic backgrounds of the normalized remainders in combinatorics and number theory on the Bernoulli numbers and polynomials, the Stirling numbers and polynomials, and their generalizations by Broder, Carlitz, and Howard, present several basic properties of normalized remainders, and collect several guesses and problems posed while investigating normalized remainders of several elementary functions.
2. Origins in mathematical analysis
We now start out to retrospect origins of inventing the concept of the normalized remainders.
2.1. First origin related to tangent
On 5 April 2023, Professor Chao-Ping Chen, a mathematician at Henan Polytechnic University, guessed that the ratio $\frac{\ln F(x)}{\ln G(x)}$ is decreasing on $\bigl(0,\frac{\pi}{2}\bigr)$, where
$F(x)=\begin{dcases}\frac{3(\tan x-x)}{x^{3}},&0<|x|<\frac{\pi}{2};\\ 1,&x=0\end{dcases}$ (2.1)
and
$G(x)=\begin{dcases}\frac{\tan x}{x},&0<|x|<\frac{\pi}{2};\\ 1,&x=0.\end{dcases}$ (2.2)
This guess was solved and applied in [54] and in the proof of [67, Theorem 3.1].
From the series expansion
$\tan x$ $=\sum_{k=1}^{\infty}\frac{2^{2k}\bigl(2^{2k}-1\bigr)}{(2k)!}|B_{2k}|x^{2k-1},\quad|x|<\frac{\pi}{2}$
$=x+\frac{x^{3}}{3}+\frac{2x^{5}}{15}+\frac{17x^{7}}{315}+\frac{62x^{9}}{2835}+\frac{1382x^{11}}{155925}+\frac{21844x^{13}}{6081075}+\dotsm$
in [16, p. 42], we observed that
$F(x)=T_{2}[\tan x]\quad\text{and}\quad G(x)=T_{0}[\tan x].$
In [32], we expanded the function $\ln T_{2}[\tan x]$ into the Maclaurin power series in two different forms. This result can be regarded as a generalization of the series expansion
$\ln T_{0}[\tan x]$ $=\sum_{k=1}^{\infty}\frac{2^{2k}\bigl(2^{2k-1}-1\bigr)}{k(2k)!}|B_{2k}|x^{2k}$
$=\frac{x^{2}}{3}+\frac{7x^{4}}{90}+\frac{62x^{6}}{2835}+\frac{127x^{8}}{18900}+\frac{146x^{10}}{66825}+\dotsm,\quad|x|<\frac{\pi}{2}$
in [16, p. 55], where $B_{2k}$ denotes the Bernoulli numbers which are generated in [68, p. 3] by
$\frac{x}{\operatorname{e}^{x}-1}=\sum_{k=0}^{\infty}B_{k}\frac{x^{k}}{k!}=1-\frac{x}{2}+\sum_{k=1}^{\infty}B_{2k}\frac{x^{2k}}{(2k)!},\quad|x|<2\pi.$ (2.3)
2.2. Second origin related to sine
In the paper [35], motivated by the functions $F(x)$ and $G(x)$ defined in (2.1) and (2.2), we considered the functions
$H(x)=\begin{dcases}\frac{6}{x^{2}}\biggl(1-\frac{\sin x}{x}\biggr),&0<|x|<\infty;\\ 1,&x=0\end{dcases}$
and
$R_{1}(x)=\begin{dcases}\frac{\ln H(x)}{\ln\frac{\sin x}{x}},&|x|\in(0,\pi);\\ \frac{3}{10},&x=0;\\ 0,&x=\pm\pi.\end{dcases}$ (2.4)
Comparing with the well-known series expansion
$\sin x$ $=\sum_{k=0}^{\infty}(-1)^{k}\frac{x^{2k+1}}{(2k+1)!},\quad x\in\mathbb{R}$
$=x-\frac{x^{3}}{6}+\frac{x^{5}}{120}-\frac{x^{7}}{5040}+\frac{x^{9}}{362880}-\dotsm,$
we observed that
$H(x)=T_{2}[\sin x]\quad\text{and}\quad\frac{\sin x}{x}=T_{0}[\sin x].$ (2.5)
In the paper [35, Theorem 1], motivated by the Maclaurin power series expansion of $\ln T_{2}[\tan x]$ obtained in [32], we expanded the logarithm $\ln T_{2}[\sin x]$ into a Maclaurin series. This result can be regarded as a generalization of the series expansion
[…]
for $|x|<\pi$ in [16, p. 55].
In the paper [35, Theorem 2], stimulated by the decreasing monotonicity obtained in [54], we proved that the ratio $R_{1}(x)$ defined by (2.4) decreasingly maps from $0,\pi]$ onto $\bigl[0,\frac{3}{10}\bigr]$.
2.3. Third origin related to cosine
In the paper [33], motivated by the above observations and results in [32, 35, 54], we discussed the functions
$I(x)=\begin{dcases}\frac{2(1-\cos x)}{x^{2}},&0<|x|<2\pi;\\ 1,&x=0\end{dcases}$
and
$R_{2}(x)=\begin{dcases}\frac{\ln I(x)}{\ln\cos x},&0<|x|<\frac{\pi}{2};\\ \frac{1}{6},&x=0;\\ 0,&x=\pm\frac{\pi}{2}.\end{dcases}$ (2.7)
Comparing with the series expansion
$\cos x$ $=\sum_{k=0}^{\infty}(-1)^{k}\frac{x^{2k}}{(2k)!}$
$=1-\frac{x^{2}}{2}+\frac{x^{4}}{24}-\frac{x^{6}}{720}+\frac{x^{8}}{40320}-\dotsm,\quad x\in\mathbb{R},$
we observed that
$I(x)=T_{1}[\cos x].$
In [33, Theorem 1], we expanded the logarithm $\ln T_{1}[\cos x]$ into a Maclaurin series. This expansion can be regarded as a generalization of the series expansion
$\ln\cos x$ $=-\sum_{k=1}^{\infty}\frac{2^{2k-1}(2^{2k}-1)}{k(2k)!}|B_{2k}|x^{2k}$ (2.8)
$=-\frac{x^{2}}{2}-\frac{x^{4}}{12}-\frac{x^{6}}{45}-\frac{17x^{8}}{2520}-\dotsm,\quad|x|<\frac{\pi}{2},$
which can be found in [16, p. 42].
In [33, Theorem 2], we proved that the ratio $R_{2}(x)$ defined by (2.7) decreasingly maps $\bigl[0,\frac{\pi}{2}\bigr]$ onto $\bigl[0,\frac{1}{6}\bigr]$.
In [33, Remark 2], Theorem 1 in [33] was simplified as
$\ln T_{1}[\cos x]=-\sum_{k=1}^{\infty}\frac{|B_{2k}|}{k}\frac{x^{2k}}{(2k)!},\quad|x|<2\pi.$ (2.9)
In [33, Remark 3], we constructed a positive and even function
$J_{n}(x)=\begin{dcases}-\frac{n+1}{|B_{2n+2}|}\frac{(2n+2)!}{x^{2n+2}}\Biggl[\ln T_{1}[\cos x]+\sum_{k=1}^{n}\frac{|B_{2k}|}{k}\frac{x^{2k}}{(2k)!}\Biggr],&0<|x|<2\pi\\ 1,&x=0\end{dcases}$ (2.10)
for $n\in\mathbb{N}_{0}$. We observed that
$J_{n}(x)=T_{2n+1}[\ln T_{1}[\cos x]],\quad n\in\mathbb{N}_{0}.$
Further, in the conclusion section [33, Section 5], we first explicitly invented the concept and used the terminology of the normalized remainder as follows.
Let $f(x)$ be an even, positive, and analytic function on (r,r)(-r,r) such that $f(0)=1$ and $f^{(2m)}(0)\neq 0$ for $m\geq 1$. Then
$f(x)=\sum_{k=0}^{\infty}f^{(2k)}(0)\frac{x^{2k}}{(2k)!},\quad x\in(-r,r).$
What are the Maclaurin power series expansions of the logarithmic expressions $\ln f(x)$ and $\ln\frac{2[f(x)-1]}{f^{\prime\prime}(0)x^{2}}$? What is the monotonicity of the even function $\frac{\ln\frac{2[f(x)-1]}{f^{\prime\prime}(0)x^{2}}}{\ln f(x)}$ on the interval $(0,r)$? Generally, what about the properties for the logarithms of the normalized remainders
$F_{n}(x)=\begin{dcases}0,&x=0\\ \ln\Biggl\{\frac{(2n)!}{f^{(2n)}(0)}\frac{1}{x^{2n}}\Biggl[f(x)-\sum_{k=0}^{n-1}f^{(2k)}(0)\frac{x^{2k}}{(2k)!}\Biggr]\Biggr\},&x\neq 0\end{dcases}$
for $n\geq 0$ and their ratios $R_{m,n}(x)=\frac{F_{n}(x)}{F_{m}(x)}$ for $n>m\geq 0$?
3. Recent results of normalized remainders
We now in a position to review main results of normalized remainders of several elementary functions.
3.1. Normalized remainders of sine and cosine functions
As a continuation of Sections 2.2 and 2.3, we considered the general normalized remainders of the cosine and sine functions in several works.
3.1.1.
In the paper [70], we considered the normalized remainder
$T_{2n-1}[\cos x]=\begin{dcases}(-1)^{n}\frac{(2n)!}{x^{2n}}\Biggl[\cos x-\sum_{k=0}^{n-1}(-1)^{k}\frac{x^{2k}}{(2k)!}\Biggr],&x\neq 0\\ 1,&x=0\end{dcases}$
for $n\in\mathbb{N}_{0}$, where an empty sum is understood to be 0.
It is easy to see that
$T_{-1}[\cos x]=\cos x\quad\text{and}\quad T_{1}[\cos x]=I(x).$
In [70, Theorem 1], we expanded the logarithm $\ln T_{2n-1}[\cos x]$ for $n\in\mathbb{N}_{0}$ into a Maclaurin series. This gives a generalization of [33, Theorem 1] and the series expansions (2.8) and (2.9).
In [70, Theorem 2], we proved the following conclusions:
  1. (1)
    For $n=0$ and $n\geq 2$, the normalized remainder $T_{2n-1}[\cos x]$ is decreasing and logarithmically concave on $\bigl(0,\frac{\pi}{2}\bigr)$.
  2. (2)
    The normalized remainder $T_{1}[\cos x]$ is decreasing on the intervals $(0,2\pi)$ and $(x_{k},2(k+1)\pi)$ for $k\in\mathbb{N}$, while it is increasing on $(2k\pi,x_{k})$ for $k\in\mathbb{N}$, where $x_{k}\in(2k\pi,2(k+1)\pi)$ for $k\in\mathbb{N}$ is the zero of the equation $\tan\frac{x}{2}=\frac{x}{2}$ on $(0,\infty)$.
In [70, Theorem 3], we verified that the ratio
$R_{3}(x)=\begin{dcases}\frac{\ln T_{3}[\cos x]}{\ln\cos x},&0<|x|<\frac{\pi}{2}\\ \frac{1}{15},&x=0\\ 0,&x=\pm\frac{\pi}{2}\end{dcases}$
is decreasing on $\bigl[0,\frac{\pi}{2}\bigr]$. This monotonicity result generalizes [33, Theorem 2], because the ratio $R_{3}(x)$ is a generalization of $R_{2}(x)$ defined in (2.7).
3.1.2.
In [81], we discussed the normalized remainder $T_{2n-1}[\cos x]$ and
$T_{2n}[\sin x]=\begin{dcases}(-1)^{n}\frac{(2n+1)!}{x^{2n+1}}\Biggl[\sin x-\sum_{k=0}^{n-1}(-1)^{k}\frac{x^{2k+1}}{(2k+1)!}\Biggr],&x\neq 0\\ 1,&x=0\end{dcases}$
for $n\in\mathbb{N}_{0}$, where an empty sum is understood to be 0.
It is obvious that those equalities in (2.5) are valid.
Among other things, the main results in [81] can be recited as follows:
  1. (1)
    The normalized remainders $T_{2n}[\sin x]$ and $T_{2n-1}[\cos x]$ for $n\in\mathbb{N}$ and $x\in\mathbb{R}$ have the integral representations
    $T_{2n}[\sin x]$ $=\begin{dcases}2n(2n+1)\int_{0}^{1}(1-u)^{2n-1}\frac{\sin(xu)}{x}\operatorname{d\!}u,&x\neq 0\\ 1,&x=0\end{dcases}$
    $=(2n+1)\int_{0}^{1}(1-u)^{2n}\cos(xu)\operatorname{d\!}u$
    and
    $T_{2n-1}[\cos x]$ $=\begin{dcases}2n(2n-1)\int_{0}^{1}(1-u)^{2n-2}\frac{\sin(xu)}{x}\operatorname{d\!}u,&x\neq 0\\ 1,&x=0\end{dcases}$
    $=2n\int_{0}^{1}(1-u)^{2n-1}\cos(xu)\operatorname{d\!}u.$
  2. (2)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n}[\sin x]$ is positive and decreasing in $x\in(0,\infty)$, while it is concave in $x\in(0,\pi)$.
  3. (3)
    The normalized remainder $T_{1}[\cos x]$ is nonnegative on $(0,\infty)$, is decreasing on $0,2\pi]$, and is concave on $(0,x_{0})$, where $x_{0}\in\bigl(\frac{\pi}{2},\pi\bigr)$ is the first positive zero of the equation
    $\bigl(x^{2}-2\bigr)\sin x+2x\cos x=0.$
    For $n\geq 2$, the normalized remainder $T_{2n-1}[\cos x]$ is positive and decreasing on $(0,\infty)$ and is concave on $(0,\pi)$.
  4. (4)
    For $n\in\mathbb{N}$, the normalized remainders $T_{2n}[\sin x]$ and $T_{2n-1}[\cos x]$ have the relations
    $T_{2n}[\sin x]={}_{1}F_{2}\biggl(1;n+1,n+\frac{3}{2};-\frac{x^{2}}{4}\biggr)$ (3.1)
    and
    $T_{2n-1}[\cos x]={}_{1}F_{2}\biggl(1;n+\frac{1}{2},n+1;-\frac{x^{2}}{4}\biggr),$ (3.2)
    where the generalized hypergeometric function $_{1}F_{2}(a;b,c;z)$ is defined in [16, p. 1020] by
    $_{1}F_{2}(a;b,c;z)=\sum_{n=0}^{\infty}\frac{(a)_{n}}{(b)_{n}(c)_{n}}\frac{z^{n}}{n!},\quad z\in\mathbb{C}$
    for $a\in\mathbb{C}$, $b,c\in\mathbb{C}\setminus\{0,-1,-2,\dotsc\}$, and
    $(z)_{n}=\prod_{\ell=0}^{n-1}(z+\ell)=\begin{cases}z(z+1)\dotsm(z+n-1),&n\in\mathbb{N};\\ 1,&n=0.\end{cases}$
    See also [30, Remark 7].
3.1.3.
In the paper [41], we considered the ratios
$\frac{T_{2n-1}[\cos x]}{T_{2n}[\sin x]},\quad\frac{T_{2n-1}[\cos x]}{T_{2n+1}[\cos x]},\quad\frac{T_{2n}[\sin x]}{T_{2n+2}[\sin x]}$
in $x\in(0,\infty)$ for $n\in\mathbb{N}$. We obtained the following conclusions:
  1. (1)
    The limits
    $\lim_{x\to 0}\frac{T_{2n-1}[\cos x]}{T_{2n}[\sin x]}$ $=1,$ $\lim_{x\to 0}\frac{T_{2n-1}[\cos x]}{T_{2n+1}[\cos x]}$ $=1,$
    $\lim_{x\to 0}\frac{T_{2n}[\sin x]}{T_{2n+2}[\sin x]}$ $=1,$ $\lim_{x\to\infty}\frac{T_{2n}[\sin x]}{T_{2n+2}[\sin x]}$ $=\frac{n(2n+1)}{(n+1)(2n+3)}$
    are valid for $n\in\mathbb{N}$, while the limits
    $\lim_{x\to\infty}\frac{T_{2n-1}[\cos x]}{T_{2n}[\sin x]}=\frac{2n-1}{2n+1}$
    and
    $\lim_{x\to\infty}\frac{T_{2n-1}[\cos x]}{T_{2n+1}[\cos x]}=\frac{n(2n-1)}{(n+1)(2n+1)}$
    are valid for $n\geq 2$.
  2. (2)
    For $n\in\mathbb{N}$, the limit
    $\lim_{x\to\infty}\frac{\int_{0}^{1}(1-u)^{2n}\cos(xu)\operatorname{d\!}u}{\int_{0}^{1}(1-u)^{2n+2}\cos(xu)\operatorname{d\!}u}=\frac{n}{n+1}$
    is valid. For $n\geq 2$, the limits
    $\lim_{x\to\infty}\frac{\int_{0}^{1}(1-u)^{2n-1}\cos(xu)\operatorname{d\!}u}{\int_{0}^{1}(1-u)^{2n}\cos(xu)\operatorname{d\!}u}=\frac{2n-1}{2n}$
    and
    $\lim_{x\to\infty}\frac{\int_{0}^{1}(1-u)^{2n-1}\cos(xu)\operatorname{d\!}u}{\int_{0}^{1}(1-u)^{2n+1}\cos(xu)\operatorname{d\!}u}=\frac{2n-1}{2n+1}$
    are valid.
  3. (3)
    The ratios $\frac{T_{1}[\cos x]}{T_{2}[\sin x]}$ and $\frac{T_{1}[\cos x]}{T_{3}[\cos x]}$ have infinitely many minimums 0 at $x=2k\pi$ for $k\in\mathbb{N}$, the ratios
    $\frac{T_{1}[\cos x]}{T_{2}[\sin x]},\quad\frac{T_{1}[\cos x]}{T_{3}[\cos x]},\quad\frac{T_{2}[\sin x]}{T_{4}[\sin x]},\quad\frac{T_{3}[\cos x]}{T_{4}[\sin x]},\quad\frac{T_{3}[\cos x]}{T_{5}[\cos x]}$
    are not monotonic in $x\in(0,\infty)$, and the ratio $\frac{T_{4}[\sin x]}{T_{6}[\sin x]}$ is decreasing in $x\in(0,\infty)$.
  4. (4)
    For $n\geq 3$, the ratios
    $\frac{T_{2n-1}[\cos x]}{T_{2n}[\sin x]},\quad\frac{T_{2n-1}[\cos x]}{T_{2n+1}[\cos x]},\quad\frac{T_{2n}[\sin x]}{T_{2n+2}[\sin x]}$
    are decreasing in $x\in(0,\infty)$.
  5. (5)
    For $x\in(0,\infty)$, the following double inequalities hold:
    $\frac{2n-1}{2n+1}$ $<\frac{T_{2n-1}[\cos x]}{T_{2n}[\sin x]}<1,\quad n\geq 3,$
    $\frac{n(2n-1)}{(n+1)(2n+1)}$ $<\frac{T_{2n-1}[\cos x]}{T_{2n+1}[\cos x]}<1,\quad n\geq 3,$
    and
    $\frac{n(2n+1)}{(n+1)(2n+3)}$ $<\frac{T_{2n}[\sin x]}{T_{2n+2}[\sin x]}<1,\quad n\geq 2.$
  6. (6)
    Using the formulas (3.1) and (3.2), as done in [30, Remark 7], we reformulated in [41, Section 5] all the above results related to the normalized remainders $T_{2n}[\sin x]$ and $T_{2n-1}[\cos x]$ in terms of the generalized hypergeometric function $_{1}F_{2}(a;b,c;z)$.
3.1.4.
In the paper [43], the authors investigated the normalized remainder
$J_{n}(x)=T_{2n+1}[\ln T_{1}[\cos x]]$
defined by (2.10) for $n\in\mathbb{N}_{0}$. They discovered the following results:
  1. (1)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{2n+1}[\ln T_{1}[\cos x]]$ is
    1. (a)
      even in $x\in(-2\pi,2\pi)$,
    2. (b)
      increasing in $x\in(0,2\pi)$,
    3. (c)
      logarithmically convex in $x\in(-2\pi,2\pi)$.
  2. (2)
    For $n\in\mathbb{N}_{0}$, the logarithm $\ln T_{2n+1}[\ln T_{1}[\cos x]]$ was expanded into a Maclaurin series.
  3. (3)
    For $n\in\mathbb{N}_{0}$, the ratio $\frac{T_{2n+3}[\ln T_{1}[\cos x]]}{T_{2n+1}[\ln T_{1}[\cos x]]}$ is increasing in $x\in(0,2\pi)$.
3.1.5.
In the paper [75], the authors considered the normalized remainder
$T_{2n+1}[\ln\sec x]=\\ \begin{dcases}\frac{(n+1)(2n+2)!\Bigl[\ln\sec x-\sum_{k=1}^{n}\frac{2^{2k-1}(2^{2k}-1)}{k(2k)!}|B_{2k}|x^{2k}\Bigr]}{2^{2n+1}(2^{2n+2}-1)|B_{2n+2}|x^{2n+2}},&x\neq 0\\ 1,&x=0\end{dcases}$
for $n\in\mathbb{N}_{0}$ and $x\in\bigl(-\frac{\pi}{2},\frac{\pi}{2}\bigr)$, where an empty sum is understood to be 0.
Since $\ln\sec x=-\ln\cos x$, it easily follows that
$T_{2n+1}[\ln\sec x]=T_{2n+1}[\ln\cos x]$ (3.3)
for $n\in\mathbb{N}_{0}$ and $x\in\bigl(-\frac{\pi}{2},\frac{\pi}{2}\bigr)$.
In the paper [75], they gained the following two conclusions:
  1. (1)
    The normalized remainder $T_{2n+1}[\ln\sec x]$ for given $n\in\mathbb{N}_{0}$ is a logarithmically convex function in $x\in\bigl(-\frac{\pi}{2},\frac{\pi}{2}\bigr)$.
  2. (2)
    The ratio $\frac{T_{2n+3}[\ln\sec x]}{T_{2n+1}[\ln\sec x]}$ for $n\in\mathbb{N}_{0}$ is an increasing function of $x\in\bigl(0,\frac{\pi}{2}\bigr)$ and a decreasing function of $x\in\bigl(-\frac{\pi}{2},0\bigr)$.
3.2. Normalized remainders of tangent function and its square
Recall from the monographs [39, Chapter XIII], [64], and [73, Chapter IV] that,
  1. (1)
    an infinitely differentiable real function $f(x)$ defined on an interval $I$ is said to be absolutely monotonic in $x\in I$ if and only if all of its derivatives satisfy $f^{(n)}(x)\geq 0$ for $n\in\mathbb{N}_{0}$ and $x\in I$;
  2. (2)
    an infinitely differentiable real function $f(x)$ defined on $I$ is said to be completely monotonic in $x\in I$ if and only if all of its derivatives satisfy $(-1)^{n}f^{(n)}(x)\geq 0$ for $n\in\mathbb{N}_{0}$ and $x\in I$.
A function $f(x)$ is completely monotonic on $(a,b)$ if and only if it is absolutely monotonic on (b,a)(-b,-a); see [73, p. 145, Definition 2c].
In [19, Definition 1] and [55, Definition 1], the concepts of logarithmically absolutely (completely) monotonic functions were defined as follows:
  1. (1)
    A positive function $f(x)$ is said to be logarithmically absolutely monotonic on an interval $I$ if it has derivatives of all orders and $\ln f(x)]^{(n)}\geq 0$ for $x\in I$ and $n\in\mathbb{N}$.
  2. (2)
    A positive function $f(x)$ is said to be logarithmically completely monotonic on an interval $I$ if it has derivatives of all orders and the inequality $(-1)^{n}[\ln f(x)]^{(n)}\geq 0$ holds for $x\in I$ and $n\in\mathbb{N}$.
In [19, Theorem 1], the authors proved that a logarithmically absolutely monotonic function on an interval $I$ is also absolutely monotonic on $I$, but not conversely. In the paper [7], [19, Theorem 4], and [55, Theorem 1], the authors proved that a logarithmically completely monotonic function on an interval $I$ is also completely monotonic on $I$, but not conversely.
3.2.1.
As a continuation of the study in Section 2.1, we considered in the paper [80] the general normalized remainder
$T_{2n-2}[\tan x]=\begin{dcases}\frac{(2n)!\bigl[\tan x-\sum_{k=1}^{n-1}\frac{2^{2k}(2^{2k}-1)}{(2k)!}|B_{2k}|x^{2k-1}\bigr]}{2^{2n}\bigl(2^{2n}-1\bigr)|B_{2n}|x^{2n-1}},&0<|x|<\frac{\pi}{2}\\ 1,&x=0\end{dcases}$
for $n\in\mathbb{N}$. Our main results are the following ones:
  1. (1)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n-2}[\tan z]$ is even on $\bigl(-\frac{\pi}{2},\frac{\pi}{2}\bigr)$, is absolutely monotonic in $z\in\bigl(0,\frac{\pi}{2}\bigr)$, and is completely monotonic in $z\in\bigl(-\frac{\pi}{2},0\bigr)$.
  2. (2)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n-2}[\tan x]$ is logarithmically absolutely monotonic in $x\in\bigl(0,\frac{\pi}{2}\bigr)$.
  3. (3)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n-2}[\tan x]$ is logarithmically completely monotonic in $x\in\bigl(-\frac{\pi}{2},0\bigr)$.
  4. (4)
    For $n\in\mathbb{N}$, the ratio $\frac{(T_{2n-2}[\tan x])^{\prime}}{xT_{2n-2}[\tan x]}$ is absolutely monotonic in $x\in\bigl(0,\frac{\pi}{2}\bigr)$.
  5. (5)
    For $n\in\mathbb{N}$, the ratio $\frac{(T_{2n-2}[\tan x])^{\prime}}{xT_{2n-2}[\tan x]}$ is completely monotonic in $x\in\bigl(-\frac{\pi}{2},0\bigr)$.
  6. (6)
    For $n\in\mathbb{N}$, the function $1-\frac{T_{2n-2}[\tan x]}{T_{2n}[\tan x]}$ is absolutely monotonic in $x\in\bigl(0,\frac{\pi}{2}\bigr)$ and is completely monotonic in $x\in\bigl(-\frac{\pi}{2},0\bigr)$. Consequently, the ratio $\frac{T_{2n-2}[\tan x]}{T_{2n}[\tan x]}$ for $n\in\mathbb{N}$ is decreasing in $x\in\bigl(0,\frac{\pi}{2}\bigr)$, is increasing in $x\in\bigl(-\frac{\pi}{2},0\bigr)$, and is concave in $x\in\bigl(-\frac{\pi}{2},\frac{\pi}{2}\bigr)$.
  7. (7)
    For $n\in\mathbb{N}$, the logarithm $\ln T_{2n-2}[\tan x]$ was expanded into a Maclaurin power series.
3.2.2.
In [76], we built the normalized remainder
$T_{2n-1}[\tan^{2}x]=\\ \begin{dcases}\frac{\displaystyle(2n+2)!\Biggl[\tan^{2}x-\sum_{\ell=1}^{n-1}\frac{2^{2\ell+2}\bigl(2^{2\ell+2}-1\bigr)(2\ell+1)}{(2\ell+2)!}|B_{2\ell+2}|x^{2\ell}\Biggr]}{2^{2n+2}(2^{2n+2}-1)(2n+1)|B_{2n+2}|x^{2n}},&x\neq 0\\ 1,&x=0\end{dcases}$
$n\in\mathbb{N}$ and $|x|<\frac{\pi}{2}$.
Due to the identity $\sec^{2}x=1+\tan^{2}x$, we derived in [76] that
$T_{2n-1}[\tan^{2}x]=T_{2n-1}[\sec^{2}x],\quad n\in\mathbb{N}.$ (3.4)
Our main results in [76] are as follows:
  1. (1)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n-1}[\tan^{2}x]$ is even, positive, and increasing on $\bigl(0,\frac{\pi}{2}\bigr)$.
  2. (2)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n-1}[\tan^{2}x]$ is logarithmically convex on $\bigl(-\frac{\pi}{2},\frac{\pi}{2}\bigr)$.
  3. (3)
    For $n\in\mathbb{N}$, the logarithm $\ln T_{2n-1}[\tan^{2}x]$ was expanded into a Maclaurin series.
3.2.3.
In [36], the ratio $\frac{T_{2n+1}[\tan^{2}x]}{T_{2n-1}[\tan^{2}x]}$ for $n\in\mathbb{N}$ was proved to be increasing in $x\in\bigl(-\frac{\pi}{2},0\bigr)$ and to be decreasing in $x\in\bigl(0,\frac{\pi}{2}\bigr)$.
3.3. Normalized remainders of exponential function
3.3.1.
In the paper [77], we considered the function $\frac{x}{\operatorname{e}^{x}-1}$ in (2.3) and investigated its normalized remainder
$T_{2n+1}\biggl[\frac{x}{\operatorname{e}^{x}-1}\biggr]=\begin{dcases}\frac{(2n+2)!}{B_{2n+2}}\frac{1}{x^{2n+2}}\Biggl[\frac{x}{\operatorname{e}^{x}-1}-1+\frac{x}{2}-\sum_{j=1}^{n}B_{2j}\frac{x^{2j}}{(2j)!}\Biggr],&x\neq 0\\ 1,&x=0\end{dcases}$
for $n\in\mathbb{N}_{0}$ and $x\in\mathbb{R}$, where an empty sum is assumed to be 0.
Our main results in [77] are recited as follows:
  1. (1)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n+1}\bigl[\frac{x}{\operatorname{e}^{x}-1}\bigr]$ is even on $(-\infty,\infty)$.
  2. (2)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n+1}\bigl[\frac{x}{\operatorname{e}^{x}-1}\bigr]$ is positive in $x\in(0,\infty)$.
  3. (3)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n+1}\bigl[\frac{x}{\operatorname{e}^{x}-1}\bigr]$ is decreasing in $x\in(0,\infty)$.
  4. (4)
    For $n\in\mathbb{N}$, the ratio $\frac{T_{2n+1}[\frac{x}{\operatorname{e}^{x}-1}]}{T_{2n+3}[\frac{x}{\operatorname{e}^{x}-1}]}$ is increasing in $x\in(0,\infty)$.
  5. (5)
    For $n\in\mathbb{N}$, the ratio $\frac{B_{2n-1}(t)}{B_{2n+1}(t)}$ is increasing in $t\in\bigl(0,\frac{1}{2}\bigr)$ and decreasing in $t\in\bigl(\frac{1}{2},1\bigr)$, where the classical Bernoulli polynomials $B_{k}(t)$ are generated in [68, p. 3] by
    $\frac{x\operatorname{e}^{xt}}{\operatorname{e}^{x}-1}=\sum_{k=0}^{\infty}B_{k}(t)\frac{x^{k}}{k!},\quad|x|<2\pi.$ (3.5)
We noticed that the generating function $\frac{x}{\operatorname{e}^{x}-1}$ of the Bernoulli numbers $B_{k}$ is just the reciprocal $\frac{1}{T_{0}[\operatorname{e}^{x}]}$ and that the generating function $\frac{x\operatorname{e}^{xt}}{\operatorname{e}^{x}-1}$ of the Bernoulli polynomials $B_{k}(t)$ is just the reciprocal $\frac{1}{T_{0}[(\operatorname{e}^{x}-1)\operatorname{e}^{-tx}]}$ or $\frac{\operatorname{e}^{xt}}{T_{0}[\operatorname{e}^{x}]}$. As a result, we derive
$T_{2n+1}\biggl[\frac{x}{\operatorname{e}^{x}-1}\biggr]=T_{2n+1}\biggl[\frac{1}{T_{0}[\operatorname{e}^{x}]}\biggr]$
for $n\in\mathbb{N}_{0}$ and $x\in\mathbb{R}$.
By [20, Theorem 2.1], we deduce that the logarithm $\ln T_{0}[\operatorname{e}^{x}]$ is convex on $(-\infty,\infty)$, is $3$-convex (that is, $\frac{\operatorname{d}^{3}\ln T_{0}[\operatorname{e}^{x}]}{\operatorname{d\!}x^{3}}\geq 0$) on $(-\infty,0)$, and is $3$-concave (that is, $\frac{\operatorname{d}^{3}\ln T_{0}[\operatorname{e}^{x}]}{\operatorname{d\!}x^{3}}\leq 0$) on $(0,\infty)$.
Motivated by [77, Proposition 1], Yang–Qi [74] and Pinelis [44] further investigated the monotonicity and other properties of the ratios
$\frac{B_{2\ell-1}(s)}{B_{2\ell+1}(s)},\quad\frac{B_{2\ell}(s)}{B_{2\ell+1}(s)},\quad\frac{B_{2m}(s)}{B_{2\ell}(s)},\quad\frac{B_{2\ell}(s)}{B_{2\ell-1}(s)},\quad\ell,m\in\mathbb{N}.$
Since 2023, when studying properties of normalized remainders of the Maclaurin power series expansions, we discovered some monotonicity results and inequalities on sequences and functions involving the ratio $\bigl|\frac{B_{2k+2}}{B_{2k}}\bigr|$ and the Riemann zeta function $\zeta(t)$. For more details, please refer to the papers [6] and [50, Section 1.12].
3.3.2.
In the paper [5], we obtained the following results:
  1. (1)
    The logarithm $\ln T_{0}[\operatorname{e}^{x}]$ was expanded into the series
    $\ln T_{0}[\operatorname{e}^{x}]$ $=\frac{x}{2}+\sum_{k=1}^{\infty}\frac{B_{2k}}{2k}\frac{x^{2k}}{(2k)!}$
    $=\frac{x}{2}+\frac{x^{2}}{24}-\frac{x^{4}}{2880}+\frac{x^{6}}{181440}-\frac{x^{8}}{9676800}+\dotsm,\quad|x|<2\pi.$
  2. (2)
    The logarithm $\ln T_{1}[\operatorname{e}^{x}]$ was expanded into the series
    $\ln T_{1}[\operatorname{e}^{x}]$ $=\frac{x}{3}+2\sum_{k=1}^{\infty}\frac{A_{k+1}}{k+1}\frac{x^{k+1}}{(k+1)!}$
    $=\frac{x}{3}+\frac{x^{2}}{36}+\frac{x^{3}}{810}-\frac{x^{4}}{12960}-\frac{x^{5}}{68040}-\frac{x^{6}}{12247200}+\frac{x^{7}}{6123600}+\dotsm,$
    where $A_{k}$ are generated (see [26, p. 979, Eq. (2.9)]) by
    $\frac{1}{T_{1}[\operatorname{e}^{x}]}$ $=\sum_{k=0}^{\infty}A_{k}\frac{x^{k}}{k!}$ (3.6)
    $=1-\frac{x}{3}+\frac{x^{2}}{36}+\frac{x^{3}}{540}-\frac{x^{4}}{6480}-\frac{x^{5}}{27216}-\frac{x^{6}}{4082400}+\dotsm$
    for $|x|<|x_{0}|$, where $x_{0}\neq 0$ is the zero, closest to the origin $x=0$, of the equation $\operatorname{e}^{x}-1-x=0$ on the complex plane $\mathbb{C}$.
    In [61, Theorem 2.1], a closed-form expression for $A_{j}$ was given by
    $A_{k}=\frac{1}{(2k)!!}\sum_{\ell=0}^{k}\frac{\binom{k}{\ell}}{\binom{k+\ell}{\ell}}\sum_{n=0}^{k-\ell}n!\langle-2\ell\rangle_{k-\ell-n}\binom{k-\ell}{n}\binom{k+\ell}{n}\\ \times\sum_{m=0}^{\ell}(-1)^{m}\binom{\ell}{m}\sum_{p=0}^{k+\ell}2^{p}\langle\ell-m\rangle_{k+\ell-p}\binom{k+\ell}{p}\frac{S(m+p,m)}{\binom{m+p}{m}}$, (3.7)
    where the falling factorial $\langle\rho\rangle_{k}$ is defined by
    $\langle\rho\rangle_{k}=\prod_{k=0}^{k-1}(\rho-k)=\begin{cases}\rho(\rho-1)\dotsm(\rho-k+1),&k\in\mathbb{N}\\ 1,&k=0\end{cases}$
    for $\rho\in\mathbb{C}$ and the Stirling numbers of the second kind $S(j,k)$ for $j\geq k\in\mathbb{N}_{0}$ can be analytically generated [13, p. 51] by
    $(T_{0}[\operatorname{e}^{x}])^{\ell}=\sum_{k=0}^{\infty}\frac{S(k+\ell,\ell)}{\binom{k+\ell}{\ell}}\frac{x^{k}}{k!},\quad\ell\in\mathbb{N}_{0}.$ (3.8)
  3. (3)
    The logarithm $\ln T_{n-1}[\operatorname{e}^{x}]$ of the normalized remainder
    $T_{n-1}[\operatorname{e}^{x}]=\begin{dcases}\frac{n!}{x^{n}}\Biggl(\operatorname{e}^{x}-\sum_{k=0}^{n-1}\frac{x^{k}}{k!}\Biggr),&x\neq 0\\ 1,&x=0\end{dcases}$
    for $n\in\mathbb{N}$ and $x\in\mathbb{R}$ was expanded into a Maclaurin series.
  4. (4)
    The Bernoulli numbers $B_{2k}$ and the Howard numbers $A_{k}$ were expressed by $B_{2k}=-2kD_{2k}(1)$ for $k\in\mathbb{N}$ and by
    $A_{k}=(-1)^{k-1}\frac{k}{2}D_{k}(2),\quad k\geq 2,$ (3.9)
    where the determinant $D_{\ell}(n)$ is defined by
    $D_{\ell+1}(n)$ $=\begin{vmatrix}\frac{1}{\binom{n+1}{1}}&\frac{\binom{0}{0}}{\binom{n}{0}}&0&0&0&\dotsm&0&0\\ \frac{1}{\binom{n+2}{2}}&\frac{\binom{1}{0}}{\binom{n+1}{1}}&\frac{\binom{1}{1}}{\binom{n}{0}}&0&0&\dotsm&0&0\\ \frac{1}{\binom{n+3}{3}}&\frac{\binom{2}{0}}{\binom{n+2}{2}}&\frac{\binom{2}{1}}{\binom{n+1}{1}}&\frac{\binom{2}{2}}{\binom{n}{0}}&0&\dotsm&0&0\\ \frac{1}{\binom{n+4}{4}}&\frac{\binom{3}{0}}{\binom{n+3}{3}}&\frac{\binom{3}{1}}{\binom{n+2}{2}}&\frac{\binom{3}{2}}{\binom{n+1}{1}}&\frac{\binom{3}{3}}{\binom{n}{0}}&\dotsm&0&0\\ \frac{1}{\binom{n+5}{5}}&\frac{\binom{4}{0}}{\binom{n+4}{4}}&\frac{\binom{4}{1}}{\binom{n+3}{3}}&\frac{\binom{4}{2}}{\binom{n+2}{2}}&\frac{\binom{4}{3}}{\binom{n+1}{1}}&\dotsm&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ \frac{1}{\binom{n+\ell-1}{\ell-1}}&\frac{\binom{\ell-2}{0}}{\binom{n+\ell-2}{\ell-2}}&\frac{\binom{\ell-2}{1}}{\binom{n+\ell-3}{\ell-3}}&\frac{\binom{\ell-2}{2}}{\binom{n+\ell-4}{\ell-4}}&\frac{\binom{\ell-2}{3}}{\binom{n+\ell-5}{\ell-5}}&\dotsm&\frac{\binom{\ell-2}{\ell-2}}{\binom{n}{0}}&0\\ \frac{1}{\binom{n+\ell}{\ell}}&\frac{\binom{\ell-1}{0}}{\binom{n+\ell-1}{\ell-1}}&\frac{\binom{\ell-1}{1}}{\binom{n+\ell-2}{\ell-2}}&\frac{\binom{\ell-1}{2}}{\binom{n+\ell-3}{\ell-3}}&\frac{\binom{\ell-1}{3}}{\binom{n+\ell-4}{\ell-4}}&\dotsm&\frac{\binom{\ell-1}{\ell-2}}{\binom{n+1}{1}}&\frac{\binom{\ell-1}{\ell-1}}{\binom{n}{0}}\\ \frac{1}{\binom{n+\ell+1}{\ell+1}}&\frac{\binom{\ell}{0}}{\binom{n+\ell}{\ell}}&\frac{\binom{\ell}{1}}{\binom{n+\ell-1}{\ell-1}}&\frac{\binom{\ell}{2}}{\binom{n+\ell-2}{\ell-2}}&\frac{\binom{\ell}{3}}{\binom{n+\ell-3}{\ell-3}}&\dotsm&\frac{\binom{\ell}{\ell-2}}{\binom{n+2}{2}}&\frac{\binom{\ell}{\ell-1}}{\binom{n+1}{1}}\end{vmatrix}$
    $=\begin{vmatrix}A_{\ell+1,1}(n)&B_{\ell+1,\ell}(n)\end{vmatrix}_{(\ell+1)\times(\ell+1)}$
    for $n\in\mathbb{N}$ and $\ell\in\mathbb{N}_{0}$, in which the matrices $A_{\ell+1,1}(n)$ and $B_{\ell+1,\ell}(n)$ are defined by
    $A_{\ell+1,1}(n)=\begin{pmatrix}\alpha_{i,j}(n)\end{pmatrix}_{1\leq i\leq\ell+1,j=1},\quad\alpha_{i,1}(n)=\frac{1}{\binom{n+i}{i}}$
    and
    $B_{\ell+1,\ell}(n)=\begin{pmatrix}\beta_{i,j}(n)\end{pmatrix}_{1\leq i\leq\ell+1,1\leq j\leq\ell},\quad\beta_{i,j}=\begin{dcases}0,&\ell\geq j>i\in\mathbb{N};\\ \frac{\binom{i-1}{j-1}}{\binom{n+i-j}{i-j}},&1\leq j\leq i\leq\ell+1.\end{dcases}$
    To some extent, the determinantal expression (3.9) is simpler and more beautiful than the formula (3.7).
  5. (5)
    For $n\in\mathbb{N}$, the ratio
    $R_{n,0}(x)=\begin{dcases}\frac{\ln T_{n-1}[\operatorname{e}^{x}]}{x},&x\neq 0\\ \frac{1}{n+1},&x=0\end{dcases}$ (3.10)
    is increasing on $(-\infty,\infty)$.
  6. (6)
    The normalized remainder $T_{n-1}[\operatorname{e}^{x}]$ is increasing and logarithmically convex on $(-\infty,\infty)$. This conclusion generalizes the first result in [20, Theorem 2.1].
  7. (7)
    For $n\in\mathbb{N}$ and $x\in\mathbb{R}$, we have
    $\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{n+j+2}{n}}\frac{x^{j}}{j!}\Biggr]\Bigg[\sum_{j=0}^{\infty}\frac{1}{\binom{n+j}{n}}\frac{x^{j}}{j!}\Biggr]\geq\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{n+j+1}{n}}\frac{x^{j}}{j!}\Biggr]^{2}.$ (3.11)
3.3.3.
In the paper [47], we gained the following results:
  1. (1)
    For $n\in\mathbb{N}$ and $x\in\mathbb{R}$, the normalized remainder $T_{n-1}[\operatorname{e}^{x}]$ is positive.
  2. (2)
    For $n\in\mathbb{N}$, the ratio $\frac{T_{n}[\operatorname{e}^{x}]}{T_{n-1}[\operatorname{e}^{x}]}$ is decreasing in $x\in(0,\infty)$.
  3. (3)
    For $n\in\mathbb{N}$, the normalized remainder $T_{n-1}[\operatorname{e}^{x}]$ is an absolutely monotonic function in $x\in\mathbb{R}$.
3.3.4.
In [78, Theorem 1], we verified the strict inequality
$\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+2}{n+1}}\frac{x^{j}}{j!}\Biggr]\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{j+n}{n}}\frac{x^{j}}{j!}\Biggr]\\ <\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n+1}}\frac{x^{j}}{j!}\Biggr]\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n}}\frac{x^{j}}{j!}\Biggr]$ (3.12)
for $x\in\mathbb{R}$ and $n\in\mathbb{N}_{0}$. In [78, Theorem 2], by virtue of the inequality (3.12), we proved that the ratio $\frac{T_{n+1}[\operatorname{e}^{x}]}{T_{n}[\operatorname{e}^{x}]}$ for $n\in\mathbb{N}_{0}$ is decreasing in $x\in\mathbb{R}$.
Theorem 2 in [78] extends [47, Theorem 1] from $x\in(0,\infty)$ to $x\in\mathbb{R}$.
3.3.5.
In the paper [31], by establishing the integral representation
$T_{n}[\operatorname{e}^{x}]=(n+1)\int_{0}^{1}v^{n}\operatorname{e}^{x(1-v)}\operatorname{d\!}v$ (3.13)
for $n\in\mathbb{N}_{0}$ and $x\in\mathbb{R}$, and making use of two monotonicity rules in [2, pp. 10–11, Theorem 1.25] and in [48, Lemma 9], [53, Lemma 12], and [58, Remark 7.2], respectively, we concisely and elegantly proved the following properties of the normalized remainder $T_{n}[\operatorname{e}^{x}]$ once again:
  1. (1)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{n}[\operatorname{e}^{x}]$ is positive in $x\in\mathbb{R}$. See also [47, Section 2].
  2. (2)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{n}[\operatorname{e}^{x}]$ is an increasing and logarithmically convex function of $x\in\mathbb{R}$. See also [5, Corollary 1].
  3. (3)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{n}[\operatorname{e}^{x}]$ is an absolutely monotonic function in $x\in\mathbb{R}$. See also [47, Theorem 2].
  4. (4)
    For $n\in\mathbb{N}$, the function $R_{n,0}(x)$ defined by (3.10) is increasing in $x\in\mathbb{R}$. See also [5, Theorem 2].
  5. (5)
    For $x\in\mathbb{R}$ and $n\in\mathbb{N}_{0}$, the inequalities (3.11) and (3.12) with equality are true. See also [5, Corollary 2] and [78, Theorem 1].
  6. (6)
    For $n\in\mathbb{N}_{0}$, the ratio $\frac{T_{n+1}[\operatorname{e}^{x}]}{T_{n}[\operatorname{e}^{x}]}$ is decreasing in $x\in\mathbb{R}$. See also [47, Theorem 1] and [78, Theorem 2].
3.4. Normalized remainders of inverse trigonometric functions
In [37], we discussed the normalized remainders
$T_{2n}[\arcsin x]=\begin{dcases}\frac{[(2n)!!]^{2}}{(2n)!}\frac{2n+1}{x^{2n+1}}\Biggl(\arcsin x-\sum_{k=0}^{n-1}\frac{(2k)!}{[(2k)!!]^{2}}\frac{x^{2k+1}}{2k+1}\Biggr),&x\neq 0\\ 1,&x=0\end{dcases}$
and
$T_{2n}[\arctan x]=\begin{dcases}(-1)^{n}\frac{2n+1}{x^{2n+1}}\Biggl[\arctan x-\sum_{k=0}^{n-1}(-1)^{k}\frac{x^{2k+1}}{2k+1}\Biggr],&x\neq 0\\ 1,&x=0\end{dcases}$
for $n\in\mathbb{N}_{0}$. We derived the following results:
  1. (1)
    For $n\in\mathbb{N}_{0}$, the normalized remainders $T_{2n}[\arcsin x]=T_{2n}[\arccos x]$ and $T_{2n}[\arctan x]$ are positive and even in $x\in(-1,1)$.
  2. (2)
    For $n\in\mathbb{N}_{0}$, the normalized remainders $T_{2n}[\arcsin x]=T_{2n}[\arccos x]$ are increasing and the normalized remainder $T_{2n}[\arctan x]$ is decreasing in $x\in(0,1)$.
  3. (3)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{2n}[\arcsin x]$ is logarithmically convex in $x\in(0,1)$.
  4. (4)
    For $n\in\mathbb{N}_{0}$ and $|x|<1$, the normalized remainder
    $T_{2n}[\operatorname{arctanh}x]=\begin{dcases}\frac{2n+1}{x^{2n+1}}\Biggl(\operatorname{arctanh}x-\sum_{k=0}^{n-1}\frac{x^{2k+1}}{2k+1}\Biggr),&x\neq 0\\ 1,&x=0\end{dcases}$
    is logarithmically convex in $x\in(0,1)$.
  5. (5)
    The functions $\ln\bigl(\frac{2}{\pi}\arccos x\bigr)$, $\ln T_{2n}[\arctan x]$, and $\ln T_{2n}[\arcsin x]$ were expanded into the Maclaurin series.
  6. (6)
    For $n\in\mathbb{N}_{0}$, the ratio $\frac{T_{2n}[\arcsin x]}{T_{2n+2}[\arcsin x]}$ is decreasing in $x\in(0,1)$.
3.5. Normalized remainders of a function involved in an integral representation of the reciprocal of the gamma function
In [68, p. 71, Eq. (3.38)], we find the integral representation
$\frac{1}{\Gamma(z)}=\frac{\operatorname{e}^{z}z^{1-z}}{\pi}\int_{0}^{\pi}\operatorname{e}^{-z\Phi(\theta)}\operatorname{d\!}\theta,\quad\Re(z)\geq 0,$
where the classical Euler gamma function $\Gamma(z)$ can be defined [68, Chapter 3] by
$\Gamma(z)=\lim_{n\to\infty}\frac{n!n^{z}}{\prod_{k=0}^{n}(z+k)},\quad z\in\mathbb{C}\setminus\{0,-1,-2,\dotsc\}$
and
$\Phi(\theta)$ $=1-\frac{\theta}{\tan\theta}+\ln\frac{\theta}{\sin\theta}$ (3.14)
$=\sum_{j=1}^{\infty}\frac{2j+1}{2j}|B_{2j}|\frac{(2\theta)^{2j}}{(2j)!}$
$=\frac{1}{2}\theta^{2}+\frac{1}{36}\theta^{4}+\frac{1}{405}\theta^{6}+\frac{1}{4200}\theta^{8}+\frac{1}{42525}\theta^{10}+\dotsm,\quad|\theta|<\pi.$
By the way, combining the series expansion (2.6) with the series expansion (3.14), we derive
$\frac{\theta}{\tan\theta}=1-\sum_{k=1}^{\infty}|B_{2k}|\frac{(2\theta)^{2k}}{(2k)!},\quad|\theta|<\pi.$
This can also be derived by the Laurent series expansion of $\cot\theta$ in the handbooks [1, p. 75, Entry 4.3.70] and [42, p. 117, Entry 4.19.6]. See also [79, Remark 1]. As a result, the function $1-\frac{\theta}{\tan\theta}$ is absolutely monotonic on $(0,\pi)$ and is completely monotonic on $(-\pi,0)$.
3.5.1.
In the paper [79], we considered the function $\Phi(\theta)$ and its normalized remainder
$T_{2n+1}[\Phi(\theta)]=\\ \begin{dcases}\frac{2n+2}{2n+3}\frac{1}{|B_{2n+2}|}\frac{(2n+2)!}{(2\theta)^{2n+2}}\Biggl[\Phi(\theta)-\sum_{j=1}^{n}\frac{2j+1}{2j}|B_{2j}|\frac{(2\theta)^{2j}}{(2j)!}\Biggr],&\theta\neq 0\\ 1,&\theta=0\end{dcases}$
for $n\in\mathbb{N}_{0}$ and $\theta\in(-\pi,\pi)$. Our main results are as follows:
  1. (1)
    The normalized remainder $T_{2n+1}[\Phi(\theta)]$ for $n\in\mathbb{N}_{0}$ is positive in $\theta\in(-\pi,\pi)$, is decreasing in $\theta\in(-\pi,0)$, and is increasing in $\theta\in(0,\pi)$.
  2. (2)
    The normalized remainder $T_{2n+1}[\Phi(\theta)]$ for $n\in\mathbb{N}_{0}$ is convex in $\theta\in(-\pi,\pi)$.
  3. (3)
    The normalized remainder $T_{2n+1}[\Phi(\theta)]$ for $n\in\mathbb{N}_{0}$ is an absolutely monotonic function in $\theta\in(0,\pi)$ and is a completely monotonic function in $\theta\in(-\pi,0)$.
  4. (4)
    The normalized remainder $T_{2n+1}[\Phi(\theta)]$ for $n\in\mathbb{N}_{0}$ is logarithmically convex in $\theta\in(-\pi,\pi)$.
  5. (5)
    The ratio $\frac{T_{2n+3}[\Phi(\theta)]}{T_{2n+1}[\Phi(\theta)]}$ for $n\in\mathbb{N}_{0}$ is decreasing in $\theta\in(-\pi,0)$ and is increasing in $\theta\in(0,\pi)$.
3.5.2.
In the paper [65], we generalized the main results in [79] and presented the following conclusions:
  1. (1)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{2n+1}[\Phi(\theta)]$ is a logarithmically absolutely monotonic function in $\theta\in(0,\pi)$ and a logarithmically completely monotonic function in $\theta\in(-\pi,0)$.
  2. (2)
    For $n\in\mathbb{N}_{0}$, the function $\frac{1}{\theta}\frac{\operatorname{d\!}\,\ln T_{2n+1}[\Phi(\theta)]}{\operatorname{d\!}\theta}$ is an absolutely monotonic function in $(0,\pi)$ and a completely monotonic function in $(-\pi,0)$.
  3. (3)
    For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{2n+1}[\Phi(\theta)]$ and $\frac{1}{\theta}\frac{\operatorname{d\!}\,\ln T_{2n+1}[\Phi(\theta)]}{\operatorname{d\!}\theta}$ can be extended analytically into the complex $z$-plane and are analytic in the disc $|z|<\pi$.
  4. (4)
    For $n\in\mathbb{N}_{0}$, the ratio $\frac{T_{2n+3}[\Phi(\theta)]}{T_{2n+1}[\Phi(\theta)]}$
    1. (a)
      is an absolutely monotonic function in $\theta\in(0,\pi)$,
    2. (b)
      is a completely monotonic function in $\theta\in(-\pi,0)$,
    3. (c)
      can be extended analytically into the complex $z$-plane and is analytic in the disc $|z|<\pi$.
  5. (5)
    The function $\frac{1}{\Phi(\theta)}$ for $0<|\theta|<\pi$ was expanded into a Laurent series.
The absolute and complete monotonicity of the ratio $\frac{T_{2n+3}[\Phi(\theta)]}{T_{2n+1}[\Phi(\theta)]}$ discovered in [65, Theorem 2] and the Laurent series of the function $\frac{1}{\Phi(\theta)}$ for $0<|\theta|<\pi$ expanded in [65, Theorem 3] imply that the function $\frac{2}{\theta^{2}}-\frac{1}{\Phi(\theta)}$ is absolutely monotonic on $(0,\pi)$ and completely monotonic on $(-\pi,0)$.
3.6. Normalized remainders of complete elliptic integrals
For complex numbers $a,b,c\in\mathbb{C}$ with $c\neq 0,-1,-2,\dotsc$, the Gauss hypergeometric function $_{2}F_{1}$ is defined [40, 46] by
$_{2}F_{1}(a,b;c;z)=\sum_{n=0}^{\infty}\frac{(a)_{n}(b)_{n}}{(c)_{n}}\frac{z^{n}}{n!},\quad|z|<1,$
where $(z)_{n}$ for $z\in\mathbb{C}$ is the Pochhammer symbol, also known as the rising factorial and the shifted factorial, defined by
$(z)_{n}=\prod_{k=0}^{n-1}(z+k)=\begin{cases}z(z+1)\dotsm(z+n-1),&n\in\mathbb{N};\\ 1,&n=0.\end{cases}$
For given $x\in[0,1]$, the complete elliptic integral of the second kind $\mathscr{E}(x)$ are defined [68, p. 315] by
$\mathscr{E}(x)=\int_{0}^{\pi/2}\sqrt{1-x^{2}\sin^{2}\theta}\,\operatorname{d\!}\theta=\int_{0}^{1}\frac{\sqrt{1-x^{2}t^{2}}\,}{\sqrt{1-t^{2}}\,}\operatorname{d\!}\theta$
and can be represented [68, p. 128] by
$\begin{dcases}\mathscr{E}(x)=\frac{\pi}{2}{\,}_{2}F_{1}\biggl(-\frac{1}{2},\frac{1}{2};1;x^{2}\biggr);\\ \mathscr{E}(0)=\frac{\pi}{2};\\ \mathscr{E}(1)=1.\end{dcases}$
Let
$\mathcal{H}(x)=\frac{1-{\,}_{2}F_{1}\bigl(-\frac{1}{2},\frac{1}{2};1;x\bigr)}{1-(1-x){\,}_{2}F_{1}\bigl(\frac{1}{2},1;\frac{3}{2};x\bigr)}=\frac{1-\frac{2}{\pi}\mathscr{E}(\sqrt{x}\,)}{1+(x-1)\frac{\operatorname{arctanh}\sqrt{x}\,}{\sqrt{x}\,}}$
and
$\mathcal{G}(x)$ $=\frac{\frac{\pi}{2}\bigl[{\,}_{2}F_{1}\bigl(-\frac{1}{2},\frac{1}{2};1;x\bigr)-1\bigr]+x{\,}_{2}F_{1}\bigl(\frac{1}{2},1;\frac{3}{2};x\bigr)}{x{\,}_{2}F_{1}(1,1;2;x)}$
$=\frac{\frac{\pi}{2}-\sqrt{x}\,\operatorname{arctanh}\sqrt{x}\,-\mathscr{E}(\sqrt{x}\,)}{\ln(1-x)}.$
3.6.1.
In the paper [71, Remarks 2 and 4], we proved that the normalized remainders $T_{n}[\mathcal{H}(x)]$ and $T_{n}[\mathcal{G}(x)]$ for $n\in\mathbb{N}_{0}$ are absolutely monotonic on $(0,1)$.
3.6.2.
In the paper [23, Theorems 3 and 4], we employed a new method to concisely prove the absolute monotonicity of the normalized remainders $T_{n}[\mathcal{H}(x)]$ and $T_{n}[\mathcal{G}(x)]$ for $n\in\mathbb{N}_{0}$ on $(0,1)$ once again.
3.7. Normalized remainder of power of inverse sine function
In [17, Theorem 1], [18, Theorem 2.1], [49, Section 4], and [51, Section 6], we discovered the Maclaurin power series expansion
$\biggl(\frac{\arcsin x}{x}\biggr)^{n}=1+\sum_{\ell=1}^{\infty}\frac{\mathcal{Q}(n,2\ell)}{\binom{n+2\ell}{2\ell}}\frac{(2x)^{2\ell}}{(2\ell)!},\quad|x|<1$ (3.15)
for $n\in\mathbb{N}$ several times by different ideas and techniques, where
$\mathcal{Q}(n,2\ell)=(-1)^{\ell}\sum_{k=0}^{2\ell}\binom{n+k-1}{n-1}s(n+2\ell-1,n+k-1)\biggl(\frac{n+2\ell-2}{2}\biggr)^{k}$ (3.16)
for $n,\ell\in\mathbb{N}$ and the Stirling numbers of the first kind $s(\ell+n,n)$ for $\ell,n\in\mathbb{N}_{0}$ can be analytically generated [38, Theorem 3.14] by
$\biggl[\frac{\ln(1+x)}{x}\biggr]^{n}=\sum_{\ell=0}^{\infty}\frac{s(\ell+n,n)}{\binom{\ell+n}{n}}\frac{x^{\ell}}{\ell!},\quad|x|<1.$ (3.17)
The power series expansion (3.15) is the unification of the Maclaurin power series
$\frac{\arcsin x}{x}$ $=\sum_{\ell=0}^{\infty}\frac{[(2\ell-1)!!]^{2}}{\binom{2\ell+1}{1}}\frac{x^{2\ell}}{(2\ell)!},$
$\biggl(\frac{\arcsin x}{x}\biggr)^{2}$ $=\sum_{\ell=0}^{\infty}\frac{[(2\ell)!!]^{2}}{\binom{2\ell+2}{2}}\frac{x^{2\ell}}{(2\ell)!},$
$\biggl(\frac{\arcsin x}{x}\biggr)^{3}$ $=\sum_{\ell=0}^{\infty}\frac{[(2\ell+1)!!]^{2}}{\binom{2\ell+3}{3}}\Biggl[\sum_{k=0}^{\ell}\frac{1}{(2k+1)^{2}}\Biggr]\frac{x^{2\ell}}{(2\ell)!},$
$\biggl(\frac{\arcsin x}{x}\biggr)^{4}$ $=\sum_{\ell=0}^{\infty}\frac{[(2\ell+2)!!]^{2}}{\binom{2\ell+4}{4}}\Biggl[\sum_{k=0}^{\ell}\frac{1}{(2k+2)^{2}}\Biggr]\frac{x^{2\ell}}{(2\ell)!},$
$\biggl(\frac{\arcsin x}{x}\biggr)^{5}$ $=\frac{1}{2}\sum_{\ell=0}^{\infty}\frac{[(2\ell+3)!!]^{2}}{\binom{2\ell+5}{5}}\Biggl[\Biggl(\sum_{k=0}^{\ell+1}\frac{1}{(2k+1)^{2}}\Biggr)^{2}-\sum_{k=0}^{\ell+1}\frac{1}{(2k+1)^{4}}\Biggr]\frac{x^{2\ell}}{(2\ell)!}$
for $|x|<1$. These series expansions can be found in the literature [3, p. 444], [17, Remark 6], [18, Remark 5.2], [50, Section 1.1], [56, Section 1], [60, Section 1], [69, pp. 279 and 331–333], [82, p. 229], and [4, 8, 66], for example.
In [49, Theorem 7], see also [59, Example 3.1], the Maclaurin series expansion (3.15) was generalized as
$\biggl(\frac{\arcsin x}{x}\biggr)^{r}=1+\sum_{\ell=1}^{\infty}\Biggl[\sum_{k=1}^{2\ell}\frac{(-r)_{k}}{(2\ell+k)!}\sum_{j=1}^{k}(-1)^{j}\binom{2\ell+k}{k-j}\mathcal{Q}(j,2\ell)\Biggr](2x)^{2\ell}$
for $r\in\mathbb{R}$ and $|x|<1$.
In the paper [50], we mainly considered the normalized remainder
$T_{2n+1}\biggl[\biggl(\frac{\arcsin x}{x}\biggr)^{q}\biggr]\\ =\begin{dcases}\frac{\binom{q+2n+2}{2n+2}}{\mathcal{Q}(q,2n+2)}\frac{(2n+2)!}{(2x)^{2n+2}}\Biggl[\biggl(\frac{\arcsin x}{x}\biggr)^{q}-1-\sum_{\ell=1}^{n}\frac{\mathcal{Q}(q,2\ell)}{\binom{q+2\ell}{2\ell}}\frac{(2x)^{2\ell}}{(2\ell)!}\Biggr],&x\neq 0\\ 1,&x=0\end{dcases}$
for $q\in\mathbb{N}$, $n\in\mathbb{N}_{0}$, and $|x|<1$, where $\mathcal{Q}(q,2\ell)$ is defined by (3.16), and acquired the following conclusions:
  1. (1)
    The normalized remainder $T_{2n+1}\bigl[\bigl(\frac{\arcsin x}{x}\bigr)^{q}\bigr]$ is an even function in $x\in(-1,1)$.
  2. (2)
    For $n,\ell\in\mathbb{N}$, the central factorial numbers of the first kind $t(n,\ell)$ satisfy
    $t(n+2\ell-1,n)=0$
    and
    $\mathcal{Q}(n,2\ell)=(-1)^{\ell}t(n+2\ell,n),$
    where the central factorial numbers of the first kind $t(n,\ell)$ are defined [63, pp. 212–217, Section 6.5] by
    $x\prod_{\ell=1}^{n-1}\biggl(x+\frac{n}{2}-\ell\biggr)=\sum_{\ell=1}^{n}t(n,\ell)x^{\ell},\quad n\in\mathbb{N}$
    and can be generated [34, p. 44] by
    $\Biggl[2\ln\Biggl(\frac{x}{2}+\sqrt{1+\biggl(\frac{x}{2}\biggr)^{2}}\,\Biggr)\Biggr]^{n}=n!\sum_{\ell=n}^{\infty}t(\ell,n)\frac{x^{\ell}}{\ell!},\quad n\in\mathbb{N}.$
    Consequently, for $n\in\mathbb{N}$, we have
    $\biggl(\frac{\arcsin x}{x}\biggr)^{n}=1+\sum_{\ell=1}^{\infty}(-1)^{\ell}\frac{t(n+2\ell,n)}{\binom{n+2\ell}{2\ell}}\frac{(2x)^{2\ell}}{(2\ell)!},\quad|x|<1.$
  3. (3)
    For $n\geq\ell\in\mathbb{N}$, the central factorial numbers of the first kind $t(n,\ell)$ can be explicitly computed by
    $t(n,\ell)=Q(\ell,n-\ell),$
    where $Q(n,k)$ is defined by
    $Q(n,k)=\sum_{\ell=0}^{k}\binom{n+\ell-1}{n-1}s(n+k-1,n+\ell-1)\biggl(\frac{n+k-2}{2}\biggr)^{\ell}$
    for $n\in\mathbb{N}$ and $k\in\mathbb{N}_{0}$, with assumptions $Q(1,1)=0$ and $Q(2,0)=1$, satisfying $Q(n,2\ell)=(-1)^{\ell}\mathcal{Q}(n,2\ell)$.
  4. (4)
    The central factorial numbers of the first kind $t(2n,2\ell)$ can be explicitly computed by
    $t(2n,2\ell)=\sum_{k=2\ell-1}^{2n-1}\binom{k}{2\ell-1}s(2n-1,k)(n-1)^{k-2\ell+1},\quad n\geq\ell\in\mathbb{N}$ (3.18)
    and the central factorial numbers of the first kind $t(2n-1,2\ell-1)$ can be explicitly computed by
    $t(2n+1,2\ell+1)=\sum_{k=2\ell}^{2n}\binom{k}{2\ell}s(2n,k)\biggl(n-\frac{1}{2}\biggr)^{k-2\ell},\quad 0\leq\ell\leq n,$
    where the factor $0^{0}$ when $n=1$ in (3.18) is assumed to be 11 and $s(n,k)$ denotes the Stirling numbers of the first kind generated in (3.17).
  5. (5)
    The central factorial numbers of the first kind $t(n,m)$ can be explicitly computed by
    $t(2n,2m)=(-1)^{n-1}\Biggl([(n-1)!]^{2}\sum_{j=m-1}^{n-1}\frac{|s(j,m-1)|}{j!}\\ +\sum_{i=1}^{n-1}(-1)^{i}\Biggl[\prod_{\ell=1}^{n-1}\bigl(i+\ell^{2}\bigr)\Biggr]\Biggl[\sum_{j=m-1}^{n-1}\binom{j}{i}\frac{|s(j,m-1)|}{j!}\Biggr]\Biggr)$
    for $n\geq m\in\mathbb{N}$ and
    $t(2n+1,2m+1)=\frac{(-1)^{n}}{4^{n-m}}\Biggl([(2n-1)!!]^{2}\sum_{j=m}^{n}\frac{|s(j,m)|}{j!}\\ +\sum_{i=1}^{n}(-1)^{i}\Biggl[\prod_{\ell=1}^{n}\bigl[i+(2\ell-1)^{2}\bigr]\Biggr]\Biggl[\sum_{j=m}^{n}\binom{j}{i}\frac{|s(j,m)|}{j!}\Biggr]\Biggr)$
    for $n\geq m\in\mathbb{N}_{0}$.
  6. (6)
    For $n\geq\ell\in\mathbb{N}$, the central factorial numbers of the first kind $t(n,\ell)$ can be computed by
    $t(2n,2\ell)$ $=(-1)^{n-\ell}\sum_{1\leq j_{1}<j_{2}<\dotsm<j_{n-\ell}\leq n-1}\prod_{m=1}^{n-\ell}(j_{m})^{2},$
    $t(2n-1,2\ell-1)$ $=(-1)^{n-\ell}\sum_{1\leq j_{1}<j_{2}<\dotsm<j_{n-\ell}\leq n-1}\prod_{m=1}^{n-\ell}\biggl(j_{m}-\frac{1}{2}\biggr)^{2},$
    and
    $t(n,\ell)=(-1)^{n-\ell}\sum_{1\leq j_{1}<j_{2}<\dotsm<j_{n-\ell}\leq n-1}\prod_{k=1}^{n-\ell}\biggl(j_{k}-\frac{n}{2}\biggr).$
    Consequently, for $n\geq\ell\in\mathbb{N}$, we have
    $(-1)^{n-\ell}t(2n,2\ell)>0\quad\text{and}\quad(-1)^{n-\ell}t(2n-1,2\ell-1)>0.$
  7. (7)
    For $n\in\mathbb{N}$, let the integer Vandermonde matrix
    $V(n)=\begin{pmatrix}1&0&0&\dotsm&0&0\\ 1&1&1&\dotsm&1&1\\ 1&2&2^{2}&\dotsm&2^{n-1}&2^{n}\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 1&n-1&(n-1)^{2}&\dotsm&(n-1)^{n-1}&(n-1)^{n}\\ 1&n&n^{2}&\dotsm&n^{n-1}&n^{n}\end{pmatrix}_{(n+1)\times(n+1)}.$
    Then its inverse matrix is
    $V^{-1}(n)=\begin{pmatrix}1&\boldsymbol{0}(n)\\ -U(n)\boldsymbol{1}^{\textup{T}}(n)&U(n)\end{pmatrix}$,
    where T stands for the transpose notation, and
    $\boldsymbol{0}(n)=(\overbrace{0,0,\dotsc,0}^{n}),\quad\boldsymbol{1}(n)=(\overbrace{1,1,\dotsc,1}^{n}),$
    $U(n)=\begin{pmatrix}u_{k,i}(n)\end{pmatrix}_{1\leq k\leq n,1\leq i\leq n}=\begin{pmatrix}\displaystyle\frac{1}{i!}\sum_{j=k}^{n}\frac{(-1)^{j-i}}{(j-i)!}s(j,k)\end{pmatrix}_{1\leq k\leq n,1\leq i\leq n}.$
  8. (8)
    For $n,\ell\in\mathbb{N}$, the quantity $\mathcal{Q}(n,2\ell)$ defined by (3.16) is positive.
  9. (9)
    For $1\leq\ell\leq n\in\mathbb{N}$, we have
    $(-1)^{\ell}\Biggl((n!)^{2}\sum_{j=\ell}^{n}\frac{|s(j,\ell)|}{j!}+\sum_{i=1}^{n}(-1)^{i}\Biggl[\sum_{j=\ell}^{n}\binom{j}{i}\frac{|s(j,\ell)|}{j!}\Biggr]\prod_{k=1}^{n}\bigl(i+k^{2}\bigr)\Biggr)>0$
    and
    $(-1)^{\ell}\Biggl([(2n-1)!!]^{2}\sum_{j=\ell}^{n}\frac{|s(j,\ell)|}{j!}\\ +\sum_{i=1}^{n}(-1)^{i}\Biggl[\sum_{j=\ell}^{n}\binom{j}{i}\frac{|s(j,\ell)|}{j!}\Biggr]\prod_{k=1}^{n}\bigl[i+(2k-1)^{2}\bigr]\Biggr)>0.$
  10. (10)
    For $1\leq q\leq 5$ and $n\in\mathbb{N}_{0}$, the normalized remainder $T_{2n+1}\bigl[\bigl(\frac{\arcsin x}{x}\bigr)^{q}\bigr]$ is positive, increasing, convex, absolutely monotonic, and logarithmically convex in $x\in[0,1)$.
  11. (11)
    The logarithm $\ln T_{2n+1}\bigl[\bigl(\frac{\arcsin x}{x}\bigr)^{q}\bigr]$ was expanded into a Maclaurin series.
  12. (12)
    The partial Bell polynomials, also known as the Bell polynomials of the second kind, are denoted and defined in [12, Definition 11.2] and [13, p. 134, Theorem A] by
    $B_{n,k}(x_{1},x_{2},\dotsc,x_{n-k+1})\\ =\sum_{\begin{subarray}{c}1\leq j\leq n-k+1,\,\ell_{j}\in\{0\}\cup\mathbb{N}\\ \sum_{j=1}^{n-k+1}j\ell_{j}=n,\,\sum_{j=1}^{n-k+1}\ell_{j}=k\end{subarray}}\frac{n!}{\prod_{j=1}^{n-k+1}\ell_{j}!}\prod_{j=1}^{n-k+1}\biggl(\frac{x_{j}}{j!}\biggr)^{\ell_{j}}.$ (3.19)
    Then the identities
    $\frac{\operatorname{B}_{2n+3,1}\bigl(1,0,1,0,9,0,\dotsc,[(2n-1)!!]^{2},0,[(2n+1)!!]^{2}\bigr)}{\operatorname{B}_{2n+1,1}\bigl(1,0,1,0,9,0,\dotsc,[(2n-3)!!]^{2},0,[(2n-1)!!]^{2}\bigr)}$ $=(1+2n)^{2},$
    $\frac{\operatorname{B}_{2n+4,2}\bigl(1,0,1,0,9,0,\dotsc,[(2n-1)!!]^{2},0,[(2n+1)!!]^{2}\bigr)}{\operatorname{B}_{2n+2,2}\bigl(1,0,1,0,9,0,\dotsc,[(2n-3)!!]^{2},0,[(2n-1)!!]^{2}\bigr)}$ $=(2+2n)^{2},$
    $\frac{\operatorname{B}_{2n+5,3}\bigl(1,0,1,0,9,0,\dotsc,[(2n-1)!!]^{2},0,[(2n+1)!!]^{2}\bigr)}{\operatorname{B}_{2n+3,3}\bigl(1,0,1,0,9,0,\dotsc,[(2n-3)!!]^{2},0,[(2n-1)!!]^{2}\bigr)}\\ =(3+2n)^{2}+\frac{1}{\sum_{k=0}^{n}\frac{1}{(2k+1)^{2}}},$
    $\frac{\operatorname{B}_{2n+6,4}\bigl(1,0,1,0,9,0,\dotsc,[(2n-1)!!]^{2},0,[(2n+1)!!]^{2}\bigr)}{\operatorname{B}_{2n+4,4}\bigl(1,0,1,0,9,0,\dotsc,[(2n-3)!!]^{2},0,[(2n-1)!!]^{2}\bigr)}\\ =(4+2n)^{2}+\frac{1}{\sum_{k=0}^{n}\frac{1}{(2k+2)^{2}}},$
    and
    $\frac{\operatorname{B}_{2n+7,5}\bigl(1,0,1,0,9,0,\dotsc,[(2n-1)!!]^{2},0,[(2n+1)!!]^{2}\bigr)}{\operatorname{B}_{2n+5,5}\bigl(1,0,1,0,9,0,\dotsc,[(2n-3)!!]^{2},0,[(2n-1)!!]^{2}\bigr)}\\ =(5+2n)^{2}\frac{\bigl[\sum_{k=0}^{n+2}\frac{1}{(2k+1)^{2}}\bigr]^{2}-\sum_{k=0}^{n+2}\frac{1}{(2k+1)^{4}}}{\bigl[\sum_{k=0}^{n+1}\frac{1}{(2k+1)^{2}}\bigr]^{2}-\sum_{k=0}^{n+1}\frac{1}{(2k+1)^{4}}}$
    are valid for $n\in\mathbb{N}$.
  13. (13)
    For $n\in\mathbb{N}$ and $1\leq\ell\leq 5$, the central factorial numbers $t(2n+\ell,\ell)$ satisfy the identities
    $\frac{t(2n+3,1)}{t(2n+1,1)}$ $=-\frac{(2n+1)^{2}}{4},$
    $\frac{t(2n+5,3)}{t(2n+3,3)}$ $=-\frac{1}{4}\Biggl[(2n+3)^{2}+\frac{1}{\sum_{k=0}^{n}\frac{1}{(2k+1)^{2}}}\Biggr],$
    $\frac{t(2n+4,2)}{t(2n+2,2)}$ $=-\frac{(2n+2)^{2}}{4},$
    $\frac{t(2n+6,4)}{t(2n+4,4)}$ $=-\frac{1}{4}\Biggl[(2n+4)^{2}+\frac{1}{\sum_{k=0}^{n}\frac{1}{(2k+2)^{2}}}\Biggr],$
    and
    $\frac{t(2n+7,5)}{t(2n+5,5)}=-\frac{(2n+5)^{2}}{4}\frac{\bigl[\sum_{k=0}^{n+2}\frac{1}{(2k+1)^{2}}\bigr]^{2}-\sum_{k=0}^{n+2}\frac{1}{(2k+1)^{4}}}{\bigl[\sum_{k=0}^{n+1}\frac{1}{(2k+1)^{2}}\bigr]^{2}-\sum_{k=0}^{n+1}\frac{1}{(2k+1)^{4}}}.$
  14. (14)
    For $m,n\in\mathbb{N}$, we have the relation
    $\frac{t(n+2m+2,n)}{t(n+2m,n)}\\ =-\frac{1}{4}\frac{\operatorname{B}_{n+2m+2,n}\bigl(1,0,1,0,9,0,\dotsc,[(2m-1)!!]^{2},0,[(2m+1)!!]^{2}\bigr)}{\operatorname{B}_{n+2m,n}\bigl(1,0,1,0,9,0,\dotsc,[(2m-3)!!]^{2},0,[(2m-1)!!]^{2}\bigr)}.$
From the Maclaurin power series expansion (3.15) and the result that the quantity $\mathcal{Q}(n,2\ell)$ for $n,\ell\in\mathbb{N}$ is positive, we conclude that the normalized remainder $T_{2n+1}\bigl[\bigl(\frac{\arcsin x}{x}\bigr)^{q}\bigr]$ for $q\in\mathbb{N}$ and $n\in\mathbb{N}_{0}$ is an absolutely monotonic function in $x\in[0,1)$ and a completely monotonic function in $x\in(-1,0]$.
4. Historic backgrounds in combinatorics and number theory
In this section, we collected some concepts and notions, which can be reformulated in terms of normalized remainders, in combinatorics and number theory.
4.1. Generating functions of Carlitz–Howard’s generalizations of Bernoulli numbers and polynomials
The equations (2.3) and (3.5) demonstrate that the Bernoulli numbers $B_{n}$ can be generated by the reciprocal $\frac{1}{T_{0}[\operatorname{e}^{x}]}$ and the Bernoulli polynomials $B_{n}(t)$ can be generated by the reciprocal $\frac{1}{T_{0}[(\operatorname{e}^{x}-1)\operatorname{e}^{-tx}]}$ or $\frac{\operatorname{e}^{xt}}{T_{0}[\operatorname{e}^{x}]}$.
The generalized Bernoulli polynomials $B_{n}^{(\sigma)}(t)$ can be generated [68, p. 4] by
$\operatorname{e}^{xt}\biggl(\frac{x}{\operatorname{e}^{x}-1}\biggr)^{\sigma}=\sum_{n=0}^{\infty}B^{(\sigma)}_{n}(t)\frac{x^{n}}{n!},\quad|z|<2\pi.$
The generating function $\operatorname{e}^{xt}\bigl(\frac{x}{\operatorname{e}^{x}-1}\bigr)^{\sigma}$ can be regarded as $\frac{\operatorname{e}^{xt}}{(T_{0}[\operatorname{e}^{x}])^{\sigma}}$.
In the thesis [24] and the papers [10, 25, 26, 28, 29], among other things, Carlitz and Howard defined the sequences and polynomials $A_{k}$, $A_{k}(t)$, $A_{n,k}$, and $A_{n,k}(t)$ by (3.6) and
$\frac{x^{2}}{2}\frac{\operatorname{e}^{xt}}{\operatorname{e}^{x}-x-1}$ $=\sum_{k=0}^{\infty}A_{k}(t)\frac{x^{k}}{k!},$ (4.1)
$\frac{x^{n}}{n!}\frac{1}{\operatorname{e}^{x}-\sum_{k=0}^{n-1}\frac{x^{k}}{k!}}$ $=\sum_{k=0}^{\infty}A_{n,k}\frac{x^{k}}{k!},$ (4.2)
$\frac{x^{n}}{n!}\frac{\operatorname{e}^{xt}}{\operatorname{e}^{x}-\sum_{k=0}^{n-1}\frac{x^{k}}{k!}}$ $=\sum_{k=0}^{\infty}A_{n,k}(t)\frac{x^{k}}{k!}$ (4.3)
for $n\in\mathbb{N}$, and they examined a lot of algebraic properties of these sequences and polynomials. It is easy to see that $A_{k}(0)=A_{k}$, $A_{1,k}=B_{k}$, $A_{2,k}=A_{k}$, $A_{n,k}(0)=A_{n,k}$, $A_{1,k}(0)=B_{k}$, $A_{2,k}(0)=A_{k}$, and $A_{1,k}(t)=B_{k}(t)$ for $k\in\mathbb{N}_{0}$. The sequence $A_{k}$ for $k\in\mathbb{N}_{0}$ was recently investigated in [5, Remark 2] and [61, Section 2]. In terms of the notation $T_{n-1}[\operatorname{e}^{x}]$, the equation (4.3) can be reformulated as
$\frac{\operatorname{e}^{xt}}{T_{n-1}[\operatorname{e}^{x}]}=\sum_{k=0}^{\infty}A_{n,k}(t)\frac{x^{k}}{k!}.$
Meanwhile, the generating functions of the sequences and polynomials $B_{k}$, $B_{k}(t)$, $A_{k}$, $A_{k}(t)$, and $A_{n,k}$ for $n\in\mathbb{N}$ in (2.3), (3.5), (3.6), (4.1), and (4.2) can be respectively reformulated as
$\frac{1}{T_{0}[\operatorname{e}^{x}]},\quad\frac{\operatorname{e}^{xt}}{T_{0}[\operatorname{e}^{x}]},\quad\frac{1}{T_{1}[\operatorname{e}^{x}]},\quad\frac{\operatorname{e}^{xt}}{T_{1}[\operatorname{e}^{x}]},\quad\frac{1}{T_{n-1}[\operatorname{e}^{x}]}.$
See also [21, Remark 4], [22, Remark 4], [31, pp. 1036–1037], and [71, Remark 2].
4.2. Generating functions of Howard’s generalization of Stirling numbers of the second kind
The equation (3.8) shows that the Stirling numbers of the second kind $S(k+\ell,\ell)$ for $k,\ell\in\mathbb{N}_{0}$ can be generated by $(T_{0}[\operatorname{e}^{x}])^{\ell}$. The equation (3.6) shows that the Howard numbers $A_{k}$ can be generated by the reciprocal $\frac{1}{T_{1}[\operatorname{e}^{x}]}$.
The $r$-associate Stirling numbers of the second kind $S_{r}(k,\ell)$ for $k\geq\ell\geq 0$ and $r\in\mathbb{N}_{0}$ are defined [27, p. 303, Eq. (1.2)] by
$\Biggl(\operatorname{e}^{x}-\sum_{k=0}^{r}\frac{x^{k}}{k!}\Biggr)^{\ell}=\Biggl(\sum_{k=r+1}^{\infty}\frac{x^{k}}{k!}\Biggr)^{\ell}=\ell!\sum_{k=(r+1)\ell}^{\infty}S_{r}(k,\ell)\frac{x^{k}}{k!};$ (4.4)
see also [61, Section 1.12]. The equation (4.4) can be reformulated as
$\begin{split}(T_{r}[\operatorname{e}^{x}])^{\ell}&=\Biggl[\frac{(r+1)!}{x^{r+1}}\Biggl(\operatorname{e}^{x}-\sum_{k=0}^{r}\frac{x^{k}}{k!}\Biggr)\Biggr]^{\ell}\\ &=\frac{\ell![(r+1)!]^{\ell}}{[(r+1)\ell]!}\sum_{k=0}^{\infty}\frac{S_{r}(k+(r+1)\ell,\ell)}{\binom{k+(r+1)\ell}{k}}\frac{x^{k}}{k!}\end{split}$ (4.5)
for $\ell,r\in\mathbb{N}_{0}$. The equation (3.8) is a special case $r=0$ of (4.5). Consequently, the integer powers $(T_{r}[\operatorname{e}^{x}])^{\ell}$ for $\ell,r\in\mathbb{N}_{0}$ can be regarded as the generating functions of the Stirling numbers of the second kind $S(k,\ell)$ and the $r$-associate Stirling numbers of the second kind $S_{r}(k,\ell)$ for $k\geq\ell\in\mathbb{N}_{0}$ and $r\in\mathbb{N}_{0}$. See also [21, Remark 4], [22, Remark 4], and [71, Remark 2].
The Stirling polynomials $S_{k}(x)$ for $k\in\mathbb{N}_{0}$ are generated [15, 57] by
$\biggl(\frac{t}{1-\operatorname{e}^{-t}}\biggr)^{x+1}=\sum_{k=0}^{\infty}S_{k}(x)\frac{t^{k}}{k!}.$
It is clear that
$S_{k}(x)=B_{k}^{(x+1)}(x+1).$
The generating function of $S_{k}(x)$ can be rewritten as
$\biggl(\frac{t}{1-\operatorname{e}^{-t}}\biggr)^{x+1}=\biggl(\frac{1}{T_{0}[\operatorname{e}^{-t}]}\biggr)^{x+1}.$
4.3. Generating functions of Howard’s generalization of Stirling numbers of the first kind
The Stirling numbers of the first kind $s(n,k)$ for $n,k\in\mathbb{N}_{0}$ can be analytically generated [38, Theorem 3.14] by (3.17). The generating function $\bigl[\frac{\ln(1+x)}{x}\bigr]^{k}$ can be regarded as the power of the normalized remainder $(T_{0}[\ln(1+x)])^{k}$.
In [27], Howard defined $s_{r}(k,\ell)$ by
$\Biggl(\ln\frac{1}{1-x}-\sum_{k=1}^{r}\frac{x^{k}}{k}\Biggr)^{\ell}=\ell!\sum_{k=(r+1)\ell}^{\infty}s_{r}(k,\ell)\frac{x^{k}}{k!},\quad r\in\mathbb{N}_{0}.$ (4.6)
It is clear that $s_{0}(k,\ell)=(-1)^{k+\ell}s(k,\ell)$.
The equation (4.6) can be reformulated as
$\Biggl[(-1)^{r}\frac{r+1}{x^{r}}\Biggl(\frac{\ln(1+x)}{x}-\sum_{k=0}^{r-1}(-1)^{k}\frac{x^{k}}{k+1}\Biggr)\Biggr]^{\ell}\\ =\frac{\ell!(r+1)^{\ell}}{[(r+1)\ell]!}\sum_{k=0}^{\infty}(-1)^{k}\frac{s_{r}(k+(r+1)\ell,\ell)}{\binom{k+(r+1)\ell}{k}}\frac{x^{k}}{k!}$ (4.7)
for $r\in\mathbb{N}_{0}$. The function
$(-1)^{r}\frac{r+1}{x^{r}}\Biggl[\frac{\ln(1+x)}{x}-\sum_{k=0}^{r-1}(-1)^{k}\frac{x^{k}}{k+1}\Biggr]$
in (4.7) is just the normalized remainder $T_{r}\bigr[\frac{\ln(1+x)}{x}\bigr]$ for $r\in\mathbb{N}_{0}$ or the normalized remainder $(T_{r}[\ln(1+x)])^{\ell}$. See also [21, Remarks 5 and 8] and [22, Remarks 5 and 8].
4.4. Generating function of Broder’s generalization of Stirling numbers of the first kind
Theorem 15 in [9] reads that the $r$-Stirling numbers of the first kind for $r\in\mathbb{N}_{0}$ have the “vertical” exponential generating function
$\sum_{k}\begin{bmatrix}k+r\\ m+r\end{bmatrix}_{r}\frac{x^{k}}{k!}=\begin{dcases}\frac{1}{m!}\biggl(\frac{1}{1-x}\biggr)^{r}\biggl(\ln\frac{1}{1-x}\biggr)^{m},&m\geq 0;\\ 0,&m<0.\end{dcases}$ (4.8)
We can rewrite (4.8) in the form
$\begin{split}\biggl(\frac{1}{1+x}\biggr)^{r}\biggl[\frac{\ln(1+x)}{x}\biggr]^{m}&=\biggl(\frac{1}{1+x}\biggr)^{r}\biggl(T_{0}\biggr[\frac{\ln(1+x)}{x}\biggr]\biggr)^{m}\\ &=\sum_{k=0}^{\infty}(-1)^{k}\frac{\left[\begin{smallmatrix}k+m+r\\ m+r\end{smallmatrix}\right]_{r}}{\binom{k+m}{m}}\frac{x^{k}}{k!},\quad r,m\in\mathbb{N}_{0}\end{split}$ (4.9)
for $|x|<1$. Taking $r=0$ in (4.7) and (4.9) and comparing with (3.17) give
$\begin{bmatrix}k\\ m\end{bmatrix}_{0}=s_{0}(k,m)=(-1)^{k+m}s(k,m),\quad k,m\in\mathbb{N}_{0}.$
See also [21, Remark 6] and [22, Remark 6].
4.5. Generating function of Broder–Carlitz’s generalization of Stirling numbers of the second kind
Theorem 16 in [9] and Eq. (3.9) in [11] states that the $r$-Stirling numbers of the second kind for $r\in\mathbb{N}_{0}$ have the exponential generating function
$\sum_{k}\left\{\begin{matrix}k+r\\ m+r\end{matrix}\right\}_{r}\frac{x^{k}}{k!}=\begin{dcases}\frac{1}{m!}\operatorname{e}^{rx}(\operatorname{e}^{x}-1)^{m},&m\geq 0;\\ 0,&m<0.\end{dcases}$ (4.10)
We can reformulate (4.10) in the form
$\operatorname{e}^{rx}(T_{0}[\operatorname{e}^{x}])^{m}=\sum_{k=0}^{\infty}\frac{\left\{\begin{smallmatrix}k+m+r\\ m+r\end{smallmatrix}\right\}_{r}}{\binom{k+m}{m}}\frac{x^{k}}{k!},\quad r,m\in\mathbb{N}_{0},\quad|x|<\infty.$ (4.11)
Taking $r=0$ in (4.5) and (4.11) and comparing with (3.8) yield
$\left\{\begin{matrix}k\\ m\end{matrix}\right\}_{0}=S_{0}(k,m)=S(k,m),\quad k,m\in\mathbb{N}_{0}.$
See also [21, Remark 7] and [22, Remark 7].
4.6. Unified treatments of generating functions of Bernoulli numbers and Stirling numbers
In the paper [52], by virtue of a determinantal formula for derivatives of the ratio between two differentiable functions, in view of the Faà di Bruno formula, and with help of several identities and closed-form formulas for partial Bell polynomials $\operatorname{B}_{n,k}$ defined by (3.19), we
  1. (1)
    established thirteen Maclaurin series expansions of the functions
    $\ln\frac{\operatorname{e}^{x}+1}{2},\quad\ln\frac{\operatorname{e}^{x}-1}{x},\quad\ln\cosh x,$
    $\ln\frac{\sinh x}{x},\quad\biggl[\frac{\ln(1+x)}{x}\biggr]^{r},\quad\biggl(\frac{\operatorname{e}^{x}-1}{x}\biggr)^{r}$
    for $r=\pm\frac{1}{2}$ or $r\in\mathbb{R}$ in terms of the Dirichlet eta function $\eta(1-2k)$, the Riemann zeta function $\zeta(1-2k)$, and the Stirling numbers of the first and second kinds $s(n,k)$ and $S(n,k)$, respectively;
  2. (2)
    presented four determinantal expressions and three recursive relations for the Bernoulli numbers $B_{2n}$;
  3. (3)
    found out three closed-form formulas for the Bernoulli numbers $B_{2n}$ and the generalized Bernoulli numbers $B_{n}^{(r)}$ in terms of the Stirling numbers of the second kind $S(n,k)$;
  4. (4)
    deduced two combinatorial identities for the Stirling numbers of the second kind $S(n,k)$;
  5. (5)
    acquired two combinatorial identities, which can be regarded as diagonal recursive relations, involving the Stirling numbers of the first and second kinds $s(n,k)$ and $S(n,k)$;
  6. (6)
    recovered an integral representation and a closed-form formula, and establish an alternative explicit and closed-form formula, for the Bernoulli numbers of the second kind $b_{n}$ in terms of the Stirling numbers of the first kind $s(n,k)$;
  7. (7)
    obtained three identities connecting the Stirling numbers of the first and second kinds $s(n,k)$ and $S(n,k)$.
The highlights of the paper [52] include the unification $\bigl(\frac{\operatorname{e}^{x}-1}{x}\bigr)^{r}$ of the generating functions for the Bernoulli numbers $B_{n}$ and the Stirling numbers of the second kind $S(n,k)$, the unification $\bigl[\frac{\ln(1+x)}{x}\bigr]^{r}$ of the generating functions for the Bernoulli numbers of the second kind $b_{n}$, which can be generated [14, 45, 62] by
$\frac{x}{\ln(1+x)}=\sum_{n=0}^{\infty}b_{n}x^{n},\quad|x|<1,$
and the Stirling numbers of the first kind $s(n,k)$, and the disclosure of the transformations between these two unifications.
4.7. Revisiting generating function of Genocchi numbers
The Genocchi numbers $G_{n}$ are generated [13, p. 49] by
$\frac{2x}{\operatorname{e}^{x}+1}=\sum_{n=1}^{\infty}G_{n}\frac{x^{n}}{n!},$
which can be reformulated as
$\frac{1}{T_{-1}[\operatorname{e}^{x}+1]}=\sum_{n=0}^{\infty}G_{n+1}\frac{x^{n}}{(n+1)!}.$
4.8. Revisiting generating functions of Euler numbers and polynomials
The Euler numbers and polynomials $E_{n}$ and $E_{n}(t)$ are defined [13, p. 48] respectively by
$\frac{2\operatorname{e}^{x}}{\operatorname{e}^{2x}+1}=\sum_{k=0}^{\infty}E_{n}\frac{x^{n}}{n!},\quad|x|<\frac{\pi}{2}\quad\text{and}\quad\frac{2\operatorname{e}^{tx}}{\operatorname{e}^{x}+1}=\sum_{k=0}^{\infty}E_{n}(t)\frac{x^{n}}{n!},\quad|x|<\pi.$
Since
$\operatorname{e}^{x}+1=2+\sum_{n=1}^{\infty}\frac{x^{n}}{n!},\quad x\in\mathbb{R},$
we can regard the generating functions $\frac{2\operatorname{e}^{x}}{\operatorname{e}^{2x}+1}$ and $\frac{2\operatorname{e}^{tx}}{\operatorname{e}^{x}+1}$ as
$\frac{\operatorname{e}^{x}}{T_{-1}[\operatorname{e}^{2x}+1]}\quad\text{and}\quad\frac{\operatorname{e}^{tx}}{T_{-1}[\operatorname{e}^{x}+1]}$
respectively. Therefore, we can generalize the Euler polynomials $E_{n}(t)$ by
$\frac{\operatorname{e}^{tx}}{(T_{n}[\operatorname{e}^{x}+1])^{\sigma}}=\frac{\operatorname{e}^{tx}}{(T_{n}[\operatorname{e}^{x}])^{\sigma}}=\sum_{k=0}^{\infty}E_{n,k}^{(\sigma)}(t)\frac{x^{k}}{k!},\quad n\in\mathbb{N}_{0}.$ (4.12)
5. Basic properties of normalized remainders
In this section, we recite the following basic properties of normalized remainders.
Theorem 5.1 ([50, Theorem 1]).
If $T_{n}[f(x),x_{0}]$ for some $n\in\mathbb{N}_{0}$ exists, then the normalized remainder $T_{n}[f(x),x_{0}]$ satisfies
$T_{n}[\alpha+f(x),x_{0}]=T_{n}[f(x),x_{0}],\quad n\in\mathbb{N}_{0},$
where $\alpha$ is a real number.
Theorem 5.2 ([37, Theorem 1]).
If $T_{n}[f(x),x_{0}]$ for some $n\in\mathbb{N}_{0}$ exists, then
$T_{n}[\beta f(x),x_{0}]=T_{n}[f(x),x_{0}],\quad n\in\mathbb{N}_{0},$
where $\beta\neq 0$ is a real number.
Theorem 5.3 ([50, Theorem 1]).
If $f^{(i)}(x_{0})=0$ for $0\leq i<k$, but $f^{(k+1)}(x_{0})\neq 0$, and the normalized remainder $T_{n}[f(x),x_{0}]$ exists for some $n\geq k$, then the normalized remainder $T_{n}[f(x),x_{0}]$ satisfies
$T_{n}[f(x),x_{0}]=T_{n-k}\biggl[\frac{f(x)}{(x-x_{0})^{k}},x_{0}\biggr],\quad n\geq k\in\mathbb{N}_{0}.$
Utilizing Theorems 5.1, 5.2, and (5.3) with the relations
$\arccos z=\frac{\pi}{2}-\arcsin z,\quad\arctan z=\frac{\pi}{2}-\operatorname{arccot}z,$
we arrive at
$T_{2n}[\arcsin x]$ $=T_{2n-1}\biggl[\frac{\arcsin x}{x}\biggr],$ $T_{2n}[\arctan x]$ $=T_{2n-1}\biggl[\frac{\arctan x}{x}\biggr],$
$T_{2n}[\arccos x]$ $=T_{2n}[\arcsin x],$ $T_{2n}[\arctan x]$ $=T_{2n}[\operatorname{arccot}x],$
and
$T_{2n}\biggl[\arccos x-\frac{\pi}{2}\biggr]$ $=T_{2n-1}\biggl[\frac{\arccos x-\frac{\pi}{2}}{x}\biggr],$
$T_{2n}\biggl[\operatorname{arccot}x-\frac{\pi}{2}\biggr]$ $=T_{2n-1}\biggl[\frac{\operatorname{arccot}x-\frac{\pi}{2}}{x}\biggr].$
See also the paper [37].
The relations (3.3) and (3.4) can also be deduced from Theorems 5.1 and 5.2, respectively.
6. Problems and guesses on normalized remainders
In this section, we collect several guesses and problems posed while investigating normalized remainders of several elementary functions.
6.1. Problems on normalized remainders
The following are six problems on concrete normalized remainders.
6.1.1. First problem
Expand the logarithm $\ln T_{2n}[\tan x]$ into a Maclaurin series.
6.1.2. Second problem
Let $L_{n}(x)=\ln T_{2n-2}[\tan x]$ for $n\in\mathbb{N}$ and let
$Q_{m,n}(x)=\begin{dcases}\frac{L_{n}(x)}{L_{m}(x)},\quad 0<|x|<\frac{\pi}{2}\\ \frac{(m+1)(2m+1)\bigl(2^{2m}-1\bigr)(2^{2n+2}-1)B_{2m}B_{2n+2}}{(n+1)(2n+1)(2^{2m+2}-1)(2^{2n}-1)B_{2m+2}B_{2n}},&x=0\end{dcases}$
for $n>m\in\mathbb{N}$. It is clear that $Q_{1,2}(x)=\frac{\ln F(x)}{\ln G(x)}$, where $F(x)$ and $G(x)$ are defined in (2.1) and (2.2).
Show that the function $Q_{m,n}(x)$ for $m>n\in\mathbb{N}$ is decreasing on $\bigl(-\frac{\pi}{2},\frac{\pi}{2}\bigr)$.
6.1.3. Third problem
Expand the logarithm $\ln T_{2n}[\sin x]$ for $n\geq 2$ into a Maclaurin series.
6.1.4. Fourth problem
Discuss the monotonicity and convexity of the ratio $\frac{\ln T_{2n+1}[\cos x]}{\ln T_{2n-1}[\cos x]}$ for $n\in\mathbb{N}_{0}$ on $\bigl[0,\frac{\pi}{2}\bigr]$.
6.1.5. Fifth problem
In [21, Remark 6] and [22, Remark 8], we proposed the problem: Investigate properties of the sequence $F(r,s,m,k)$ generated by
$\biggl(\frac{1}{1+x}\biggr)^{r}(T_{s}[\ln(1+x)])^{m}$ $=\biggl(\frac{1}{1+x}\biggr)^{r}\biggl(T_{s}\biggr[\frac{\ln(1+x)}{x}\biggr]\biggr)^{m}$
$=\sum_{k=0}^{\infty}F(r,s,m,k)\frac{x^{k}}{k!}$
for $r,m\in\mathbb{C}$, $s\in\mathbb{N}_{0}$, and $|x|<1$.
The integral representation (3.13) and the ideas with techniques in [52, Section 7] should be useful when studying this problem.
6.1.6. Sixth problem
In [21, Remark 7] and [22, Remark 7], we proposed the problem: Generally, we can consider a new kind of polynomials $A_{n,k}^{(\sigma)}(t)$ generated by
$\frac{\operatorname{e}^{xt}}{(T_{n-1}[\operatorname{e}^{x}])^{\sigma}}=\sum_{k=0}^{\infty}A_{n,k}^{(\sigma)}(t)\frac{x^{k}}{k!},\quad n\in\mathbb{N},\quad\sigma,x,t\in\mathbb{C}.$
It is obvious that the polynomial sequence $A_{n,k}^{(\sigma)}(t)$ is a unified generalization of $B_{k}$, $B_{k}(t)$, $A_{k}$, $A_{k}(t)$, $A_{n,k}(t)$, $A_{n,k}(t)$, and the Stirling numbers of the second kind $S(n,k)$, and the Stirling polynomials $S_{k}(x)$.
It is not difficult to see that the sequences $A_{n,k}^{(\sigma)}(t)$ and $E_{n,k}^{(\sigma)}(t)$, generated in (4.12), are essentially the same one.
The integral representation (3.13) should be useful when studying this problem.
6.2. Guesses on normalized remainders
The following are nine guesses on concrete normalized remainders.
6.2.1. First guess
In [36, Remark 4.1], we proposed the guess: The ratio $\frac{T_{2n+1}[\tan^{2}x]}{T_{2n-1}[\tan^{2}x]}$ for $n\in\mathbb{N}$ is concave in $x\in\bigl(-\frac{\pi}{2},\frac{\pi}{2}\bigr)$.
6.2.2. Second guess
In the paper [65, Section 4.1], we posed the following guesses: For $n\in\mathbb{N}$, the ratio $\frac{T_{2n-1}[\tan^{2}x]}{T_{2n+1}[\tan^{2}x]}$ is completely monotonic in $x\in\bigl(-\frac{\pi}{2},0\bigr)$ and is absolutely monotonic in $x\in\bigl(0,\frac{\pi}{2}\bigr)$.
6.2.3. Third guess
In the paper [65, Section 4.2], we posed the following guesses: For $n\in\mathbb{N}_{0}$, the normalized remainder $T_{n}[\operatorname{e}^{x}]$ is a logarithmically absolutely monotonic function in $x\in\mathbb{R}$.
6.2.4. Fourth guess
In the paper [65, Section 4.2], we posed the following guesses: For $n\in\mathbb{N}_{0}$, the ratio $\frac{T_{n}[\operatorname{e}^{x}]}{T_{n+1}[\operatorname{e}^{x}]}$ is an absolutely monotonic function in $x\in\mathbb{R}$.
6.2.5. Fifth guess
In [47, Remark 7], we guessed that the functions $G_{n}(x)$ and $H_{n}(x)$ for $n\in\mathbb{N}$ are absolutely monotonic in $x\in\mathbb{R}$, where
$G_{n}(x)=\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n+1}}\frac{x^{j}}{j!}\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n}}\frac{x^{j}}{j!}-\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+2}{n+1}}\frac{x^{j}}{j!}\sum_{j=0}^{\infty}\frac{1}{\binom{j+n}{n}}\frac{x^{j}}{j!}$
and
$H_{n}(x)=\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+2}{n}}\frac{x^{j}}{j!}\sum_{j=0}^{\infty}\frac{1}{\binom{j+n}{n}}\frac{x^{j}}{j!}-\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n}}\frac{x^{j}}{j!}\Biggr]^{2}.$
This guess is stronger than the inequalities (3.11) and (3.12).
6.2.6. Sixth guess
Since the quantity $\mathcal{Q}(q,2\ell)$ for $q,\ell\in\mathbb{N}$ is positive, we guessed in [50, Section 1] that the normalized remainder $T_{2n+1}\bigl[\bigl(\frac{\arcsin x}{x}\bigr)^{q}\bigr]$ for $q\in\mathbb{N}$ and $n\in\mathbb{N}_{0}$ is logarithmically absolutely monotonic in $x\in[0,1)$.
6.2.7. Seventh guess
In the paper [37], we guessed that the normalized remainder $T_{2n}[\arctan x]$ for $n\in\mathbb{N}_{0}$ is logarithmically concave on $(0,1)$.
6.2.8. Eighth guess
In the paper [37], we guessed that the ratio $\frac{T_{2n}[\arcsin x]}{T_{2n+2}[\arcsin x]}$ for $n\in\mathbb{N}_{0}$ is concave in $x\in(-1,1)$.
6.2.9. Ninth guess
In the paper [37], we guessed that the ratio $\frac{T_{2n}[\arctan x]}{T_{2n+2}[\arctan x]}$ for $n\in\mathbb{N}_{0}$ is increasing in $x\in(0,\infty)$.
7. Conclusions
Normalized remainders represent a novel concept in mathematical analysis and combinatorial number theory. This concept is grounded in well-established mathematical knowledge and supported by a rich historical foundation. The main results reviewed in this work serve as initial steps and a fundamental basis for further exploration. These findings illustrate that the study of normalized remainders is both intriguing and significant, with promising prospects for future development.